### v1

# Dissipative qutrit-mediated stable charging

Abstract
In this work, we propose a stable charging scheme mediated by a dissipative three-level system, which allows a unidirectional energy flow from the external power source to a finite-dimensional quantum battery. By virtue of the dissipation, the battery avoids the spontaneous discharging induced by the time-reversal symmetry in any unitary charging scheme. Irrespective to the initial state, the battery can be eventually stabilized at the maximal-ergotropy state as long as the charger-battery interaction is present. We use a Dyson series of Lindbladian superoperator to obtain an effective master equation to describe the battery dynamics and extract the optimization condition for charging efficiency. Also it proves that the Fermi’s golden rule can adapt to the non-Hermitian Hamiltonian. We justify the outstanding performance of our charging scheme when applied to the finite-size system with a uniform energy ladder operator, the large spin battery, and the truncated harmonic oscillator battery.
I Introduction
As a microscopic unit for energy storage, quantum battery Campaioli et al. (2024a) has recently received remarkable attention as both a promising energy supply device for quantum technology Auffèves (2022) and a tool to develop the wide application of quantum thermodynamics Horodecki and Oppenheim (2013); Hovhannisyan et al. (2013); Vinjanampathy and Anders (2016); Campaioli et al. (2018). Since formally introduced by Alicki and Fannes Alicki and Fannes (2013), a lot of efforts have been devoted to studying the role of quantum resources such as quantum correlation and quantum coherence in the charging performance of various quantum batteries Alicki and Fannes (2013); Campaioli et al. (2017); Kamin et al. (2020); Andolina et al. (2019); Seah et al. (2021); Arjmandi et al. (2022a); Shi et al. (2022); Yang et al. (2024). It is found that the collective charging schemes could achieve a super-extensive scaling in power, which supports that the quantum correlation provides a dramatic boost to the charging efficiency of the parallel charging schemes Ferraro et al. (2018); Le et al. (2018); Rossini et al. (2020). Quantum speed up in charging can also be found in a single high-dimensional quantum battery via repeated collisions between chargers and battery Seah et al. (2021); Salvia et al. (2023); Beleño et al. (2024). In addition, quantum measurements on chargers can give rise to a more efficient charging performance than the measurement-free schemes and do not invoke the initial quantum coherence in both battery and chargers Yan and Jing (2023).
Numerous prototypes of quantum batteries have been developed in diverse models. The one-dimensional spin chain battery has been proposed in Ref. Le et al. (2018). Then many investigations further extended the spin chain to the spin-ensemble systems Qi and Jing (2021); Huangfu and Jing (2021); Zhao et al. (2021); Arjmandi et al. (2022b); Barra et al. (2022); Gao et al. (2022); Ghosh and Sen (De); Yao and Shao (2022). For example, by mediating between the charger and battery both consisting of a spin ensemble through a magnon system, we can realize a long-range charging Qi and Jing (2021). Dick model battery is another type of extensively studied model Ferraro et al. (2018); Andolina et al. (2019); Zhang et al. (2019); Crescente et al. (2020a, b), which is based on the interaction between an ensemble of noninteracting two-level systems (TLSs) and a common cavity mode. The connection and distinction between the spin-charger and cavity-charger schemes were discussed in Ref. Huangfu and Jing (2021). In addition to the spin systems, the bosonic quantum battery based on an ensemble of harmonic oscillators (HOs) was proposed in Ref. Friis and Huber (2018) focusing on the charging precision, i.e., the variance of high-level population distribution.
Besides the charging efficiency, stabilizing a quantum battery at the charged (high-ergotropy) state is also crucial for the reliable operation over quantum batteries. Many conventional schemes have to immediately decouple the charger from the battery of a finite size when the latter reaches the fully charged state. If they remain coupled with each other, energy can flow back from the battery to the charger, degrading the charging-efficiency and the energy retention Ferraro et al. (2018); Zhang et al. (2019); Andolina et al. (2018, 2019); Guo et al. (2024). Extra operations have been designed to address this challenge. The Zeno-protection scheme applies sequential projective measurements on the high energy state of a two-level quantum battery Gherardini et al. (2020). The time-dependent Hamiltonian schemes use the stimulated Raman adiabatic passage to maintain a stable charged state Santos et al. (2019); Rojo-Francàs et al. (2024). Nonlocal and strong chaotic correlation was demonstrated in a Sachdev-Ye-Kitaev quantum battery Rosa et al. (2020) to improve the charging stability, to name a few.
An alternative approach for stable charging is assisted by a dissipative charger connecting the battery and the external power source Farina et al. (2019). In that model that both charger and battery are identical HOs or TLSs, the energy and ergotropy of the battery can be stabilized at a finite value determined by the charger-battery coupling strength and the intensity of the external power source. However, the battery could not be stabilized at its highest energy level in such a setting. Introducing a common dissipative environment for the composite system of battery and charger and carefully designing their inner coupling strength and the dissipative rates can increase the steady energy in the battery Ahmadi et al. (2024).
In this work, we introduce an efficient and stable charging model that the charger is modeled as a dissipative $\Delta$-type qutrit mediating between the external power source and the battery. The charger is coupled to the battery through the energy exchange interaction. By deriving an effective master equation for the battery using a perturbation technique for Lindbliadian superoperator, we find that a unidirectional energy flow from the external source to the battery is constructed due to the time-reversal symmetry breaking by the dissipative charger. We first consider the battery with a uniform energy ladder operator that is popular in many worksSeah et al. (2021); Chen et al. (2022); Mitchison et al. (2021). Our scheme is characterized with universality in enhancing and holding the battery energy for any initial state Sathe and Caravelli (2024) and does not require the charger to be decoupled from the battery after the charging is completed. We apply our scheme to more realistic scenarios by modeling quantum battery as a large spin and a truncated quantum harmonic oscillator. It is interesting to find that choosing an appropriate charger-battery coupling strength or adjusting it during the charging process can significantly boost the charging efficiency.
The rest of the paper is organized as follows. In Sec. II, we introduce our dissipative-mediated charging model. In Sec. III, we derive an effective master equation for the battery in a weak-driving regime (see appendix A for details) and obtain the optimization condition for charging efficiency. In a broader perspective, we actually develop an efficient technique to extend the Fermi’s Golden rule to adapt to the non-Hermitian Hamiltonian. In Sec. IV, we discuss the large spin quantum battery. It is found that charging efficiency can be significantly boosted by optimizing the effective transition rate for special subspaces. In Sec. V, we discuss the truncated harmonic oscillator quantum battery. It is shown that by adjusting the charger-battery coupling strength according to the instantaneous state of the battery, the charging efficiency can be greatly enhanced. The conclusion is provided in Sec. VI.
II Model
Figure 1: Sketch of the source-charger-battery system. In our charging scheme, the dissipative $\Delta$-type qutrit serves as the charger, that is driven by an external power source and transfers energy to the battery.
Our indirect charging model consists of a quantum battery of a finite size and a dissipative charger that is driven by an external power source. As shown in Fig. 1, the charger is a dissipative $\Delta$-type qutrit. An external field directly drives the transition between levels $|g\rangle$ and $|h\rangle$. The quantum battery in Secs. II and III has $N+1$ uniform energy levels with a bare Hamiltonian ($\hbar=1$) $H_{B}=E\sum_{n=0}^{N}n|n\rangle\langle n|$. The charging process is switched on when the quantum battery couples to the qutrit via the transition $|e\rangle\leftrightarrow|h\rangle$. The full system Hamiltonian reads,
$H=$ $\omega_{hg}|h\rangle\langle h|+\omega_{eg}|e\rangle\langle e|+H_{B}$ (1)
$+\left(\Omega|h\rangle\langle g|e^{-i\omega_{L}t}+gA^{\dagger}|e\rangle\langle h|+{\rm H.c.}\right),$
where $\omega_{ij}$ with $i,j=g,e,h$ is the level splitting between qutrit states $|i\rangle$ and $|j\rangle$. $\Omega$ and $\omega_{L}$ are the intensity and the frequency of the external field, respectively. $g$ is the coupling strength between the charger and the battery. $A^{\dagger}=\sum_{n=0}^{N-1}A_{n}|n+1\rangle\langle n|$ is the creation operator of the battery, which satisfies $A^{\dagger}|N\rangle=A|0\rangle=0$ and $A_{n}$ is a dimensionless transition coefficient between $|n\rangle$ and $|n+1\rangle$. $A_{n\geq N}=0$. In the rotating frame with respect to
$U=\exp\left\{i\left[H_{B}+\omega_{L}|h\rangle\langle h|+(\omega_{L}-E)|e\rangle\langle e|\right]t\right\},$ (2)
the full Hamiltonian reads
$H^{\prime}=$ $UHU^{\dagger}-iU\frac{\partial}{\partial t}U^{\dagger}=\Delta|h\rangle\langle h|+\delta|e\rangle\langle e|$ (3)
$+\Omega\left(|h\rangle\langle g|+|g\rangle\langle h|\right)+g\left(A^{\dagger}|e\rangle\langle h|+A|h\rangle\langle e|\right)$
with detunings $\Delta=\omega_{hg}-\omega_{L}$ and $\delta=\omega_{eg}+E-\omega_{L}$. When the charger is subject to a local dissipative environment, the evolution of the charger-battery system is governed by the master equation Carmichael (1999):
$\dot{\rho}(t)=$ $-i[H^{\prime},\rho(t)]+\sum_{ij=hg,he,eg}\mathcal{L}[L_{ij}]\rho(t),$ (4)
where $L_{ij}=\sqrt{\gamma_{ij}}|j\rangle\langle i|$ is the Lindblad operator (quantum jump operator) from the qutrit level $|i\rangle$ to $|j\rangle$ with a decay rate $\gamma_{ij}$. $\mathcal{L}[o]$ is the Lindblad superoperator defined as $\mathcal{L}[o]\rho(t)\equiv o\rho(t)o^{\dagger}-1/2\{o^{\dagger}o,\rho(t)\}$.
III Effective master equation
Figure 2: (a) Main mechanism of our charging scheme. The external power source $\Omega$, the charger-battery interaction $gA_{n}$, and the spontaneous decay $\gamma_{eg}$ induces the transitions $|gn\rangle\rightarrow|hn\rangle$, $|hn\rangle\rightarrow|e(n+1)\rangle$, and $|e(n+1)\rangle\rightarrow|g(n+1)\rangle$, respectively. They give rise to an effective transition $|gn\rangle\rightarrow|g(n+1)\rangle$ with a rate $\gamma_{\rm eff}$. (b) Leading-order transitions induced by the atomic spontaneous emissions from $|h\rangle$ to $|e\rangle$ and from $|e\rangle$ to $|g\rangle$. The decay channel indicated by $\gamma_{he}$ is the main negative factor for charging.
In our charging scheme, the mediator qutrit is initially prepared in its ground state $|g\rangle$. In the weak-driving regime $\Omega\ll\Delta,\ \gamma_{hg}$, one can suppose that the qutrit subject to a dissipative environment remains mostly in its ground state $\rho(t)\approx P_{g}\rho(t)P_{g}$, where $P_{g}\equiv|g\rangle\langle g|$. According to the master equation (4), the main mechanism underlying the indirect charging can be illustrated in the subspaces spanned by $\{|gn\rangle,|hn\rangle,|e(n+1)\rangle,|g(n+1)\rangle\}$ for $n<N$. As shown in Fig. 2(a), the system is driven from $|gn\rangle$ to $|hn\rangle$ through the external power source $\Omega$ and from $|hn\rangle$ to $|e(n+1)\rangle$ via the charger-battery interaction with a coupling strength $gA_{n}$. Subsequently, the spontaneous decay $\gamma_{eg}$ induces a transition from $|e(n+1)\rangle$ to $|g(n+1)\rangle$. These processes constitute an effective unidirectional transition from $|gn\rangle$ to $|g(n+1)\rangle$, which can be quantitatively described by the effective Lindblad operator in Eq. (48):
$L_{eg,n}^{\rm eff}=\frac{\sqrt{\gamma_{eg}}\Omega gA_{n}}{g^{2}A_{n}^{2}-\tilde{\Delta}\tilde{\delta}}|g(n+1)\rangle\langle gn|$ (5)
with $\tilde{\Delta}\equiv\Delta-i\gamma_{hg}/2-i\gamma_{he}/2$ and $\tilde{\delta}\equiv\delta-i\gamma_{eg}/2$ the complex detunings. Due to the large-detuning condition, the decay from $|hn\rangle$ to $|gn\rangle$ contributes to a pure dephasing of the state $|gn\rangle$ in the leading order. The relevant effective Lindblad operator given by Eq. (47) reads
$L_{hg,n}^{\rm eff}=\frac{\sqrt{\gamma_{hg}}\Omega\tilde{\delta}}{\tilde{\Delta}\tilde{\delta}-g^{2}A_{n}^{2}}|gn\rangle\langle gn|.$ (6)
In addition, the energy shift of $|gn\rangle$ induced by the weak driving and the charger-battery interaction can be described by the effective Hamiltonian
$H_{n}^{\rm eff}=-\Omega^{2}{\rm Re}\left(\frac{\tilde{\delta}}{\tilde{\Delta}\tilde{\delta}-g^{2}A_{n}^{2}}\right)|gn\rangle\langle gn|.$ (7)
The relevant details can be found in Appendix A.
By tracing out the qutrit from Eq. (4), the effective equation of motion of the battery is given by
$\dot{\rho}_{B}(t)=-i[H_{\rm eff}^{B},\rho_{B}(t)]+\mathcal{L}[L^{B}_{hg,{\rm eff}}]\rho_{B}(t)+\mathcal{L}[L^{B}_{eg,{\rm eff}}]\rho_{B}(t),$ (8)
where
$H^{B}_{\rm eff}$ $=\sum_{n=0}^{N}-\Omega^{2}{\rm Re}\left(\frac{\tilde{\delta}}{\tilde{\Delta}\tilde{\delta}-g^{2}A_{n}^{2}}\right)|n\rangle\langle n|$ (9)
$=-\Omega^{2}{\rm Re}\left(\frac{\tilde{\delta}}{\tilde{\Delta}\tilde{\delta}-AA^{\dagger}g^{2}}\right),$
and
$L^{B}_{hg,{\rm eff}}$ $=\sum_{n=0}^{N}\frac{\sqrt{\gamma_{hg}}\Omega\tilde{\delta}}{\tilde{\Delta}\tilde{\delta}-g^{2}A_{n}^{2}}|n\rangle\langle n|=\frac{\sqrt{\gamma_{hg}}\Omega\tilde{\delta}}{\tilde{\Delta}\tilde{\delta}-AA^{\dagger}g^{2}},$ (10a)
$L^{B}_{eg,{\rm eff}}$ $=\sum_{n=0}^{N}\frac{\sqrt{\gamma_{eg}}\Omega gA_{n}}{g^{2}A_{n}^{2}-\tilde{\Delta}\tilde{\delta}}|n+1\rangle\langle n|=\frac{\sqrt{\gamma_{eg}}\Omega}{A^{\dagger}Ag^{2}-\tilde{\Delta}\tilde{\delta}}A^{\dagger}.$ (10b)
are the effective Hamiltonian and Lindblad operators in the battery Hilbert space, respectively. When expanding the coefficients of the effective Lindblad operators in terms of series
$\langle gn|L_{hg,n}^{\rm eff}|gn\rangle$ $=\sqrt{\gamma_{hg}}\sum_{k=0}^{\infty}\left(\frac{g^{2}A_{n}^{2}}{\tilde{\Delta}\tilde{\delta}}\right)^{k}\frac{\Omega}{\tilde{\Delta}}$ (11)
$=-\sqrt{\gamma_{hg}}\sum_{k=0}^{\infty}\Omega_{\rm eff}^{(2k)},$
$\langle g(n+1)|L_{eg,n}^{\rm eff}|gn\rangle$ $=-\sqrt{\gamma_{eg}}\sum_{k=0}^{\infty}\left(\frac{g^{2}A_{n}^{2}}{\tilde{\Delta}\tilde{\delta}}\right)^{k}\frac{gA_{n}}{\tilde{\delta}}\frac{\Omega}{\tilde{\Delta}}$
$=-\sqrt{\gamma_{eg}}\sum_{k=0}^{\infty}\Omega_{\rm eff}^{(2k+1)},$
it is interesting to find that $\Omega_{\rm eff}^{(l)}$’s are precisely the effective coupling strengths connecting $|i\rangle=|gn\rangle$ and $|f\rangle=|hn\rangle$ or $|e(n+1)\rangle$ up to the first order in $\Omega$ and $l$ order in $gA_{n}$ with respect to the non-Hermitian Hamiltonian in Eq. (31):
$\tilde{H}_{n}$ $=\tilde{H}_{0,n}+\tilde{V}_{n},$ (12)
$\tilde{H}_{0,n}$ $=\tilde{\Delta}|hn\rangle\langle hn|+\tilde{\delta}|e(n+1)\rangle\langle e(n+1)|,$
$\tilde{V}_{n}$ $=gA_{n}(|hn\rangle\langle e(n+1)|+|e(n+1)\rangle\langle hn|).$
Particularly, they can be obtained by the constrained high-order Fermi’s golden rules Kockum et al. (2017) with the form
$\Omega^{(l)}_{\rm eff}$ (13)
$=$ $\sum_{m_{1},m_{2},\cdots,m_{l}}\frac{\tilde{V}_{n,fm_{l-1}}\cdots\tilde{V}_{n,m_{2}m_{1}}\tilde{V}_{n,m_{1}i}}{(\tilde{E}_{i}-\tilde{E}_{m_{1}})(\tilde{E}_{i}-\tilde{E}_{m_{2}})\cdots(\tilde{E}_{i}-\tilde{E}_{m_{l-1}})},$
where $\tilde{V}_{n,m_{i}m_{j}}=\langle m_{i}|\tilde{V}_{n}|m_{j}\rangle$ with $m_{i}=hn,e(n+1)$ the intermediate states. $\tilde{E}_{m_{i}}=\langle m_{i}|\tilde{H}_{0,n}|m_{i}\rangle$ and $\tilde{E}_{i}=\langle gn|\tilde{H}_{0,n}|gn\rangle=0$ are the diagonal elements of $\tilde{H}_{0,n}$. For an even $l$, $\Omega_{\rm eff}^{(l)}$ represents the effective coupling strength for connecting $|gn\rangle$ and $|hn\rangle$ through the process $|gn\rangle\rightarrow|hn\rangle\rightarrow|e(n+1)\rangle\cdots\rightarrow|hn\rangle$. For an odd $l$, it represents the effective coupling strength for connecting $|gn\rangle$ and $|e(n+1)\rangle$ through the process $|gn\rangle\rightarrow|hn\rangle\rightarrow|e(n+1)\rangle\cdots\rightarrow|e(n+1)\rangle$. Similarly, the coefficient of the effective Hamiltonian (9) is the real part of the summation of all round trips about $|gn\rangle$.
The effective transition rate about $|gn\rangle\rightarrow|g(n+1)\rangle$ in Fig. 2(a) is given by
$\gamma_{n,{\rm eff}}$ $=|\langle g(n+1)|L_{eg,n}^{\rm eff}|gn\rangle|^{2}=\left|\frac{\sqrt{\gamma_{eg}}\Omega gA_{n}}{g^{2}A_{n}^{2}-\tilde{\Delta}\tilde{\delta}}\right|^{2}$ (14)
$=\frac{\gamma_{eg}\Omega^{2}}{g^{2}A_{n}^{2}+|\tilde{\Delta}\tilde{\delta}|^{2}/(g^{2}A_{n}^{2})-2|\tilde{\Delta}\tilde{\delta}|\cos\phi},$
where $\phi=\arg(\tilde{\Delta}\tilde{\delta})$. With fixed detunings $\tilde{\Delta}$ and $\tilde{\delta}$, one can find that the effective transition rate in the $n$th subspace can be optimized by setting
$g^{2}_{\rm opt}A^{2}_{n}=|\tilde{\Delta}\tilde{\delta}|.$ (15)
The decay channel $|h\rangle\rightarrow|e\rangle$ plays a negative role in the charging efficiency through a higher-order process than $|e\rangle\rightarrow|g\rangle$ and $|h\rangle\rightarrow|g\rangle$ (see Appendix A). As the leading-order contribution depicted in Fig. 2(b), the system in the subspace is driven by the external power from $|gn\rangle$ to $|hn\rangle$ then decay from $|hn\rangle$ to $|en\rangle$ and then decay from $|en\rangle$ to $|gn\rangle$ will give rise to an effective dephasing on the state $|gn\rangle$ with a rate given by Eq. (58). In a biased fluxonium system at about $0.05\leq\Phi_{\rm ex}/\Phi_{0}\leq 0.15$, where $\Phi_{\rm ex}$ and $\Phi_{0}$ are the applied flux and flux quantum, respectively, $\gamma_{he}$ is the same order in magnitude as $\gamma_{eg}$ and is one order of magnitude smaller than $\gamma_{hg}$ Joo et al. (2010); Manucharyan et al. (2009), i.e. $\gamma_{eg}\approx\gamma_{he}\ll\gamma_{hg}$. With such a parametric setting, one can confirm that the dephasing rate caused by the process in Fig. 2(b) is negligible in comparison to that described by Eq. (6) for $|gn\rangle\rightarrow|hn\rangle\rightarrow|gn\rangle$ in Fig. 2(a). Similarly, the process that from $|gn\rangle$ to $|hn\rangle$ through the external driving, from $|hn\rangle$ to $|en\rangle$ through the spontaneous decay $\gamma_{he}$, from $|en\rangle$ to $|h(n-1)\rangle$ through the charger-battery interaction, and finally from $|h(n-1)\rangle$ to $|g(n-1)\rangle$ through the spontaneous decay $\gamma_{hg}$ gives rise to an effective transition from $|gn\rangle$ to $|g(n-1)\rangle$ with a rate given by Eq. (61). It is also negligible comparing to the effective transition rate in Eq. (14) under the condition $|\tilde{\Delta}|\gg|\tilde{\delta}|$ and $\gamma_{eg}\approx\gamma_{he}$. Consequently, the presence of a decay channel $|h\rangle\rightarrow|e\rangle$ does not significantly alter our unidirectional energy flow.
Figure 3: Comparison between the dynamics obtained by the full master equation (4) and the effective master equation (8) of (a) battery energy $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ with various $\Omega$. The coupling strength $g$ satisfies the optimized condition (15). The battery size is set as $N=50$. $\gamma_{he}=0$, $\gamma_{hg}=10\gamma_{eg}=0.1E_{B}$, and $\Delta=10\delta=0.1E_{B}$.
Figure 4: Charging performance in the presence of the decay from $|h\rangle$ to $|e\rangle$ with $\gamma_{he}=\gamma_{eg}$ under various inverse temperature $\beta$ ($k_{B}\equiv 1$). (a) Energy stored in the battery $\Delta E(t)$. (b) Ergotropy $\mathcal{E}(t)$. The driving intensity is fixed as $\Omega=0.05\gamma_{hg}$. Other parameters are the same as in Fig. 3.
The justification of our charging scheme begins with a simple case with $A_{n}=1$, resulting in uniform effective transition rates across all the subspaces. Substituting the optimized condition (15) to Eq. (14), it is found that the optimized effective transition rate reads
$\gamma^{\rm opt}_{\rm eff}=\frac{\gamma_{eg}\Omega^{2}}{4|\tilde{\Delta}\tilde{\delta}|\sin^{2}(\phi/2)}.$ (16)
The performance of the charging protocol can be evaluated by the energy stored in the battery
$\Delta E(t)\equiv{\rm Tr}[H_{B}\rho_{B}(t)]-{\rm Tr}[H_{B}\rho_{B}(0)],$ (17)
and the ergotropy Allahverdyan et al. (2004), the maximum amount of work that can be extracted from the battery via unitary transformation, which is defined as
$\mathcal{E}(t)\equiv{\rm Tr}[H_{B}(\rho_{B}(t))]-{\rm Tr}[H_{B}\sigma(t)],$ (18)
Here $\sigma(t)$ is the passive state that is diagonal in the battery eigenstates basis with non-increasing eigenvalues of $\rho_{B}(t)$:
$\sigma(t)=\sum_{n=0}^{N}\lambda_{n}(t)|n\rangle\langle n|,\quad\lambda_{n+1}(t)\leq\lambda_{n}(t)$ (19)
Both energy and ergotropy saturate when the battery state approaches the maximal-ergotropy state $\rho_{B}\rightarrow|N\rangle\langle N|$. By virtue of the last term in Eq. (8) interpreted by Eq. (10b), our charging scheme can hold on that state without decoupling the charger from the battery after the charging process is completed.
In Fig. 3, we compare the dynamics of the battery energy $\Delta E(t)$ and ergotropy $\mathcal{E}(t)$ under the full master equation (4) and the effective master equation (8). The battery is initially prepared at its ground state $\rho_{B}(0)=|0\rangle\langle 0|$ with vanishing energy and ergotropy. It is found that for either energy in Fig. 3(a) or ergotropy in Fig. 3(b), the charging efficiency becomes higher under a stronger driving intensity $\Omega$. On the other hand, the deviation between the effective dynamics and the full dynamics increases when $\Omega$ becomes larger and larger. It is due to the breakdown of the weak driving assumption.
A practical charging scheme can transform a high-energy yet zero-ergotropy state, e.g., the thermal state, to be a high-energy and high-ergotropy state. In Fig. 4(a) and 4(b), we demonstrate the charging performance for the initial thermal states with various inverse temperatures $\beta$ by the stored energy $\Delta E(t)$ and ergotropy $\mathcal{E}(t)$, respectively. $\beta=\infty$ means the ground state. It can be found that both energy and ergotropy still monotonically increase with the charging time and ultimately stabilise with their respective maximum values. Note high-ergotropy state is different from high-energy state in population distribution. Due to the unidirectional transition from the lower Fock states to the higher ones as indicated in Eq. (10b), an initially higher-temperature battery state is more convenient to be transformed to be a higher-ergotropy state than the lower-temperature state. The final increment of $\Delta E$ for an initially high-temperature state is significantly smaller than that for $\mathcal{E}(t)$. Moveover, we take account of a nonvanishing $\gamma_{he}$ in Fig. 4. In comparison to Fig. 3, one can find that the effect of the process shown in Fig. 2(b) is marginal in the charging efficiency, that is consistent with our analysis.
IV Large spin battery
Figure 5: Effective transition rate $\gamma_{m,{\rm eff}}$ (in unit of $E_{B}$) from $|J,m\rangle$ to $|J,m+1\rangle$ in the parameter space of $g$ (in unit of $E_{B}$) and $m$. The black dashed line indicates the optimization condition in Eq. (23). The horizontal red dot-dashed line indicates the optimization condition for $m=0$. The energy quantum number is set as $J=25$. The driving intensity is $\Omega=0.1\gamma_{hg}$. The other parameters are the same as in Fig. 3.
Figure 6: Full and effective dynamics of (a) battery energy $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ under the optimization conditions of $g=g_{m=-J,{\rm opt}}$ (solid lines and dashed lines) and $g=g_{m=0,{\rm opt}}$ (dotted lines and dot-dashed line), as defined in Eq. (23). The quantum number of energy is set as $J=25$ and the decay rate is $\gamma_{he}=\gamma_{eg}$. $\Omega=0.1\gamma_{hg}$. The other parameters are the same as in Fig. 3.
To implement our charging scheme in a more physically relevant scenario, we consider an $N=2J+1$ dimensional large-spin battery in this section. The battery Hamiltonian reads $H_{B}=E_{B}J_{z}=E_{B}\sum_{m=-J}^{J}m|J,m\rangle\langle J,m|$, where $J$ is an integer or half-integer quantum number of energy. The charger and the battery interact with each other through
$H_{I}=g(J_{+}|e\rangle\langle h|+J_{-}|h\rangle\langle e|),$ (20)
where the creation operator of battery is the collective angular momentum operator
$J_{+}=\sum_{m=-J}^{J}\sqrt{J(J+1)-m(m+1)}|J,m+1\rangle\langle J,m|,$ (21)
and $J_{-}=J_{+}^{\dagger}$. The effective transition rate from $|J,m\rangle$ to $|J,m+1\rangle$ can be obtained by substituting $A_{m}=\sqrt{J(J+1)-m(m+1)}$ to Eq. (14). We have
$\gamma_{m,{\rm eff}}=\left|\frac{\sqrt{\gamma_{eg}}\Omega g\sqrt{J(J+1)-m(m+1)}}{g^{2}[J(J+1)-m(m+1)]-\tilde{\Delta}\tilde{\delta}}\right|^{2}.$ (22)
In contrast to the uniform model with $A_{n}=1$, the effective transition rates in the large spin model vary across different subspaces, the same is true for the optimization condition:
$g_{m,{\rm opt}}^{2}[J(J+1)-m(m+1)]=|\tilde{\Delta}\tilde{\delta}|.$ (23)
Figure 7: Full and effective dynamics of (a) energy stored in the battery $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ under initial thermal states with different inverse temperature $\beta$. The other parameters are the same as in Fig. 6.
In Fig. 5, we present the dependence of effective transition rate $\gamma_{m,{\rm eff}}$ on the coupling strength between the qutrit charger and the large-spin battery $g$ and the quantum number of angular momentum $m$ to confirm the optimization condition in Eq. (23), as indicated by the black dashed line. It is noted that by setting $m=0$ in Eq. (23), the relevant $\gamma_{m,{\rm eff}}$ closely approximates the optimal conditions across a substantial number of subspaces as shown by the red dash-dotted line. In practice, it is a $m$-independent setting for enhancing the charging efficiency.
In Fig. 6(a) and 6(b), we compare the charging efficiency under the optimized condition of $g=g_{m=-J,{\rm opt}}$ with that under $g=g_{m=0,{\rm opt}}$ when the battery is initialised as the ground state $\rho_{B}(0)=|J,-J\rangle\langle J,-J|$. It is found that while the optimized condition $g=g_{-J,{\rm opt}}$ exhibits an advantage at the beginning of charging due to the initial population distribution, the charging rate under $g=g_{0,{\rm opt}}$ dramatically increases as the charging proceeds. The latter quickly surpasses the former. This behavior is consistent with our effective master equation. Nevertheless, the effective dynamics under $g=g_{0,{\rm opt}}$ exhibits a moderate deviation from the full one, and it perfectly matches with the full dynamics under $g=g_{-J,{\rm opt}}$.
Now we check the case that the large-spin battery is initially prepared at the thermal states with various inverse temperature $\beta$ by the dynamics of stored energy $\Delta E(t)$ and ergotropy $\mathcal{E}(t)$ in Fig. 7(a) and 7(b), respectively. The optimization condition is chosen as $g=g_{0,{\rm opt}}$. It is found that the ergotropy can be maintained about $\mathcal{E}_{\rm steady}>0.995\mathcal{E}_{\rm max}$, irrespective to the temperature $\beta$ of the initial state, where $\mathcal{E}_{\rm max}=2JE_{B}$. Moreover, it is interesting to find that a higher initial temperature of the battery results in a higher growth rate of the ergotropy.
V Harmonic oscillator battery
Figure 8: Effective transition rate $\gamma_{n,{\rm eff}}$ (in unit of $E_{B}$) from $|n\rangle$ to $|n+1\rangle$ in the parametric space of $g$ (in unit of $E_{B}$) and $n$. The black dashed line indicates the optimization condition in Eq. (25). The battery size is chosen as $N=50$. $\Omega=0.1\gamma_{hg}$. The other parameters are the same as in Fig. 3.
In this section, we consider a quantum battery consisting of an $N+1$-dimensional truncated harmonic oscillator. In this case, the ladder operator becomes a truncated bosonic creation operator $A^{\dagger}=\sum_{n=0}^{N-1}\sqrt{n+1}|n+1\rangle\langle n|$. The effective transition rate from $|n\rangle$ to $|n+1\rangle$ can be obtained by substituting $A_{n}=\sqrt{n+1}$ to Eq. (14):
$\gamma_{n,{\rm eff}}=\left|\frac{\sqrt{\gamma_{eg}}\Omega g\sqrt{n+1}}{g^{2}(n+1)-\tilde{\Delta}\tilde{\delta}}\right|^{2}.$ (24)
Subsequently, it can be optimized under the condition
$g_{n,{\rm opt}}^{2}(n+1)=|\tilde{\Delta}\tilde{\delta}|.$ (25)
Figure 9: Full and effective dynamics of (a) battery energy $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ with an invariant coupling strength $g=g_{0,{\rm opt}}$ in Eq. (25) during the charging process (black solid lines and blue dashed lines) and with a coupling strength quenched at $\gamma_{eg}\tau$ (green and yellow lines) according to Eq. (26). The decay rate $\gamma_{he}=\gamma_{eg}$. $\Omega=0.1\gamma_{hg}$. The other parameters are the same as in Fig. 3.
The effective transition rate in the parameter space of charger-battery coupling strength $g$ and $n$ is plotted in Fig. 8. The optimization condition is indicated by the black dashed line. Distinct from the large-spin battery, it is hardly to find a subspace-independent coupling strength $g_{\rm opt}$ to optimize the effective transition rates for the harmonic-oscillator battery, since $g_{n,{\rm opt}}$ decreases with $n$ in a monotonic way. Nevertheless, the charging efficiency can be enhanced by adjusting the coupling strength during the charging process. By our scheme, one can initially optimize the effective transition rate from $|0\rangle$ to $|1\rangle$ by choosing $g=g_{0,{\rm opt}}$ in Eq. (25) due to the fact that the ground state is always the most populated state for a thermal state. Then after an interval of $\tau$, one can modify the coupling strength to be
$g=g_{\bar{n}(\tau),{\rm opt}}=\sqrt{\frac{|\tilde{\Delta}\tilde{\delta}|}{\bar{n}(\tau)+1}},$ (26)
where $\bar{n}(\tau)={\rm Tr}[\rho_{B}(\tau)H_{B}]/E_{B}$ is the instantaneous average excitation of the truncated HO battery.
We compare the charging performance with a uniform $g$ during the whole charging evolution and that with a quenched $g$ by the dynamics of $\Delta E(t)$ and $\mathcal{E}(t)$ in Figs. 9(a) and 9(b), respectively. The yellow and green lines indicate the quenched moment at $\gamma_{eg}\tau=500$ and $\gamma_{eg}\tau=1000$, respectively, in which the charger-battery coupling strength is only modified at the specified moment and calculated according to Eq. (26). The battery is initially prepared in the ground state $\rho_{B}(0)=|0\rangle\langle 0|$. It is clearly that the strategy of a uniform $g$ is the lowest in charging efficiency and both energy and ergotropy have an abrupt and significant enhancement when $g$ can be re-optimized. The numerical simulation indicates that the full dynamics is perfectly matched with the effective dynamics and the charging efficiency is slightly overestimated after the adjustment of the coupling strength $g$.
Figure 10: Full and effective dynamics of (a) energy stored in the battery $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ under initial thermal states with various $\beta$. The coupling strength is only modified at $\gamma_{eg}\tau=500$ according to Eq. (26). The other parameters are the same as in Fig. 9.
The dynamics of the stored energy $\Delta E(t)$ and ergotropy $\mathcal{E}(t)$ for charging the HO battery initially in the thermal states with a finite temperature is demonstrated in Figs. 10(a) and 10(b), respectively. The results are similar to the large-spin battery presented in Fig. 7. A higher initial temperature or energy results in a higher growth rate of the ergotropy. In addition, the steady value of the ergotropy is insensitive to the initial inverse temperature $\beta$. In this case, we have $\mathcal{E}_{\rm steady}>0.994\mathcal{E}_{\rm max}$ with $\mathcal{E}_{\rm max}=NE_{B}$.
VI Conclusion
In summary, we have proposed a stable charging scheme assisted by a dissipative $\Delta$-type qutrit, which relies on the nonreciprocal interaction between the external power source and the battery and the ensued unidirectional energy flow. Three types of finite dimensional quantum batteries, including the uniform coupling strength model, the large-spin battery, and the truncated harmonic oscillator battery, have confirmed that the battery energy and ergotropy can be maintained at their respective maximum values and the battery need not to be decoupled from the charger after the charging process is completed. By deriving the effective master equation of the battery, we can obtain the optimization condition of the charging efficiency. For the large spin battery, the charging efficiency can be significantly enhanced by choosing the charger-battery coupling strength according to the optimization condition in the subspace with $m=0$. For the truncated harmonic oscillator battery, the enhancement can be achieved by adjusting the coupling strength according to the instantaneous state of the battery during charging. In application, our work provides an efficient initial-state-independent charging scheme and avoids the discharging induced by the time-reversal symmetry. In theory, it extends the high-order Fermi’s Golden rule to the regime of non-Hermitian Hamiltonian.
Appendix A The effective master equation
This appendix contributes to deriving the effective Hamiltonian and Lindblad operators for the quantum battery $\rho_{B}(t)$ using a technique extending the James’ method James and Jerke (2007); Gamel and James (2010); Shao et al. (2017) from Hamiltonian to Lindbladian.
By defining a non-Hermitian Hamiltonian
$H_{\rm NH}$ $=H^{\prime}-\frac{i}{2}\sum_{k=hg,eg,he}L^{\dagger}_{k}L_{k}=\tilde{H}+V,$ (27)
$\tilde{H}$ $=\tilde{\Delta}|h\rangle\langle h|+\tilde{\delta}|e\rangle\langle e|+g(A^{\dagger}|e\rangle\langle h|+{\rm H.c.}),$ (28)
$V$ $=\Omega|h\rangle\langle g|+{\rm H.c.},$ (29)
where $L_{hg}$, $L_{eg}$, and $L_{he}$ are $\sqrt{\gamma_{hg}}|g\rangle\langle h|$, $\sqrt{\gamma_{eg}}|g\rangle\langle e|$, and $\sqrt{\gamma_{he}}|h\rangle\langle e|$, respectively, and $\tilde{\Delta}=\Delta-i\gamma_{hg}/2-i\gamma_{he}/2$ and $\tilde{\delta}=\delta-i\gamma_{eg}/2$ are complex detunings, the full master equation (4) can be rewritten as
$\dot{\rho}(t)=$ $-i[(\tilde{H}+V)\rho(t)-\rho(t)(\tilde{H}^{\dagger}+V)]$ (30)
$+L_{hg}\rho(t)L_{hg}^{\dagger}+L_{eg}\rho(t)L_{eg}^{\dagger}+L_{he}\rho(t)L_{he}^{\dagger}.$
In the weak-driving regime Reiter and Sørensen (2012), the subspace spanned by $\{|gn\rangle,|hn\rangle,|e(n+1)\rangle,|g(n+1)\rangle\}$ with $n<N$ in Fig. 2(a) can be divided into the rapidly-decaying exited subspace spanned by $\{|hn\rangle,|e(n+1)\rangle\}$ and the ground subspace spanned by $\{|gn\rangle,|g(n+1)\rangle\}$. With respect to Fig. 2(a), the non-Hermitian Hamiltonian defined in Eq. (28) is given by
$\tilde{H}_{n}=$ $\tilde{\Delta}|hn\rangle\langle hn|+\tilde{\delta}|e(n+1)\rangle\langle e(n+1)|$ (31)
$+gA_{n}\left[|hn\rangle\langle e(n+1)|+|e(n+1)\rangle\langle hn|\right],$
involving only with the excited subspaces. The weak external driving $V$ leads to the excitation process from the ground subspaces to the excited subspaces with a Hermitian Hamiltonian
$V_{n}=V_{+,n}+V^{\dagger}_{+,n}=\Omega|hn\rangle\langle gn|+{\rm H.c.}.$ (32)
The Lindblad operators induced by the qutrit decay read
$L_{hg,n}=\sqrt{\gamma_{hg}}|gn\rangle\langle hn|,$ (33)
$L_{eg,n}=\sqrt{\gamma_{eg}}|g(n+1)\rangle\langle e(n+1)|.$
To employ the perturbation theory, it is convenient to transform to the rotating frame with respect to $\tilde{U}(t)=e^{i\tilde{H}t}$. The master equation (30) then becomes
$\dot{\tilde{\rho}}(t)=$ $-i[\tilde{V}(t)\tilde{\rho}(t)-\tilde{\rho}(t)\tilde{V}^{\dagger}(t)]+\tilde{L}_{hg}(t)\tilde{\rho}(t)\tilde{L}^{\dagger}_{hg}(t)$ (34)
$+\tilde{L}_{eg}(t)\tilde{\rho}(t)\tilde{L}^{\dagger}_{eg}(t)+\tilde{L}_{he}(t)\tilde{\rho}(t)\tilde{L}^{\dagger}_{he}(t),$
where $\tilde{\rho}(t)=\tilde{U}(t)\rho(t)\tilde{U}^{\dagger}(t)$ is the density matrix in the rotating frame. The other operators are transformed accordingly with $\tilde{\mathcal{O}}(t)=\tilde{U}(t)\mathcal{O}\tilde{U}^{-1}(t)$, $\mathcal{O}=V$, $L_{hg}$, $L_{eg}$, and $L_{he}$.
With respect to the main process in Fig. 2(a), the transformed master equation (34) in the subspace can be written as
$\dot{\tilde{\rho}}_{n}(t)=$ $-i\left[\tilde{V}_{n}(t)\tilde{\rho}_{n}(t)-\tilde{\rho}_{n}(t)\tilde{V}_{n}^{\dagger}\right]$ (35)
$+\tilde{L}_{hg,n}(t)\tilde{\rho}_{n}(t)\tilde{L}^{\dagger}_{hg,n}(t)$
$+\tilde{L}_{eg,n}(t)\tilde{\rho}_{n}(t)\tilde{L}^{\dagger}_{eg,n}(t),$
where
$\tilde{V}_{n}(t)$ $=\Omega(e^{i\tilde{H}_{n}t}|hn\rangle\langle gn|+|gn\rangle\langle hn|e^{-i\tilde{H}_{n}t}),$ (36)
$\tilde{L}_{hg,n}(t)$ $=\sqrt{\gamma_{hg}}|gn\rangle\langle hn|e^{-i\tilde{H}_{n}t},$ (37)
$\tilde{L}_{eg,n}(t)$ $=\sqrt{\gamma_{eg}}|g(n+1)\rangle\langle e(n+1)|e^{-i\tilde{H}_{n}t}.$ (38)
The last term about $\tilde{L}_{he}(t)$ in Eq. (34) that appears only as the leading-order contribution depicted in Fig. 2(b) will be analyzed later.
Then we vectorize the density matrix of each subspace to be a $4^{2}\times 1$ vector denoted by $|\tilde{\rho}_{n}(t)\rangle\rangle$. The equation of motion in this representation is given by
$\partial_{t}|\tilde{\rho}_{n}(t)\rangle\rangle=\mathbf{L}_{n}(t)|\tilde{\rho}_{n}(t)\rangle\rangle.$ (39)
Her, $\mathbf{L}(t)$ is the $4^{2}\times 4^{2}$ superoperator in the form of
$\mathbf{L}_{n}(t)=\mathcal{V}_{n}(t)+\mathcal{J}_{n}(t),$ (40)
$\mathcal{V}_{n}(t)=-i(\mathbb{I}\otimes\tilde{V}_{n}(t)-\tilde{V}_{n}^{*}(t)\otimes\mathbb{I}),$
$\mathcal{J}_{n}(t)=\tilde{L}^{*}_{hg,n}(t)\otimes\tilde{L}_{hg,n}(t)+\tilde{L}^{*}_{eg,n}(t)\otimes\tilde{L}_{eg,n}(t),$
where $\mathbb{I}$ is the $4\times 4$ identity operator in the subspace Campaioli et al. (2024b). $\mathcal{V}(t)$ and $\mathcal{J}(t)$ represent the non-Hermitian evolution part and the quantum jump part, respectively. The formal solution of Eq. (39) is
$|\tilde{\rho}_{n}(t)\rangle\rangle=|\tilde{\rho}_{n}(0)\rangle\rangle+\int_{0}^{t}dt_{1}\mathbf{L}_{n}(t_{1})|\tilde{\rho}_{n}(t_{1})\rangle\rangle.$ (41)
In the large detuning condition, i.e., $\Delta+\delta\gg\Omega$, the transformation operator $U(t)=e^{i\tilde{H}_{n}(t)}\propto\exp[i(\tilde{\Delta}+\tilde{\delta})t/2]$ is highly oscillating with time. By iteratively substituting Eq. (41) into Eq. (39), neglecting the fast oscillating terms $\mathcal{V}(t)|\tilde{\rho}_{n}(0)\rangle\rangle$, and taking the Markovian approximation Shao et al. (2017), the equation of motion can written into a Dyson series:
$\partial_{t}|\tilde{\rho}_{n}(t)\rangle\rangle=$ $\left[\mathbf{L}_{n}(t)\int_{0}^{t}dt_{1}\mathbf{L}_{n}(t_{1})+\mathbf{L}_{n}(t)\int_{0}^{t}dt_{1}\mathbf{L}_{n}(t_{1})\int_{0}^{t_{1}}dt_{2}\mathbf{L}_{n}(t_{2})+\cdots\right]|\tilde{\rho}_{n}(t)\rangle\rangle.$ (42)
$=$ $\left[\mathbf{L}^{(2)}_{{\rm eff},n}(t)+\mathbf{L}^{(3)}_{{\rm eff},n}(t)+\cdots\right]|\tilde{\rho}_{n}(t)\rangle\rangle.$
It is assumed that the qutrit remains mostly in its ground state during the charging process $|\tilde{\rho}(t)\rangle\rangle\approx\mathcal{P}_{g}|\tilde{\rho}(t)\rangle\rangle$, where $\mathcal{P}_{g}=P_{g}\otimes P_{g}$. The equation of motion for the ground-state branch then takes the form
$\partial_{t}\mathcal{P}_{g}|\tilde{\rho}_{n}(t)\rangle\rangle\approx\mathcal{P}_{g}\left[\mathbf{L}^{(2)}_{{\rm eff},n}(t)+\mathbf{L}^{(3)}_{{\rm eff},n}(t)\right]\mathcal{P}_{g}|\tilde{\rho}_{n}(t)\rangle\rangle$ (43)
$=\mathcal{P}_{g}\bigg{[}\mathcal{V}_{n}(t)\int_{0}^{t}dt_{1}\mathcal{V}_{n}(t_{1})$
$+\mathcal{J}_{n}(t)\int_{0}^{t}dt_{1}\mathcal{V}_{n}(t_{1})\int_{0}^{t_{1}}dt_{2}\mathcal{V}_{n}(t_{2})\bigg{]}\mathcal{P}_{g}|\tilde{\rho}_{n}(t)\rangle\rangle.$
up to the second order in driving intensity $\Omega$. By substituting Eq. (36) into the first term of the Eq. (43), we can obtain the effective superoperator for a non-Hermitian evolution:
[…]
$=$ $-\left[P_{g}\otimes V_{+,n}^{\dagger}(i\tilde{H}_{n})^{-1}V_{+,n}+V_{+,n}^{\dagger}(-i\tilde{H}_{n}^{*})^{-1}V_{+,n}\otimes P_{g}\right],$
[…]
$\mathcal{H}^{\rm eff}_{n}=-V_{+,n}^{\dagger}(\tilde{H}_{n})^{-1}V_{+,n}=-\frac{\Omega^{2}\tilde{\delta}}{\tilde{\Delta}\tilde{\delta}-g^{2}A_{n}^{2}}|gn\rangle\langle gn|.$ (45)
By substituting Eqs. (36)-(38) into the second term of Eq. (43), we can obtain the effective superoperator for quantum jumps:
[…]
$=$ $L_{hg,n}(\tilde{H}_{n}^{*})^{-1}V_{+,n}\otimes L_{hg,n}(\tilde{H}_{n})^{-1}V_{+,n}$
$+L_{eg,n}(\tilde{H}_{n}^{*})^{-1}V_{+,n}\otimes L_{eg,n}(\tilde{H}_{n})^{-1}V_{+,n}.$
By transforming back to the original representation, the second-order effective Lindblad operators turn out to be
$L_{hg,n}^{\text{eff}}=$ $L_{hg,n}(\tilde{H}_{n})^{-1}V_{+,n}=\frac{\sqrt{\gamma_{hg}}\Omega\tilde{\delta}}{\tilde{\Delta}\tilde{\delta}-g^{2}A_{n}^{2}}|gn\rangle\langle gn|,$ (47)
$L_{eg,n}^{\text{eff}}=$ $L_{eg,n}(\tilde{H}_{n})^{-1}V_{+,n}=\frac{\sqrt{\gamma_{eg}}\Omega gA_{n}}{g^{2}A_{n}^{2}-\tilde{\Delta}\tilde{\delta}}|g(n+1)\rangle\langle gn|.$ (48)
Therefore according to Eq. (43), the equation of motion in the ground state subspaces reads
$P_{g}\dot{\rho}_{n}(t)P_{g}=-i\left[\mathcal{H}^{\rm eff}_{n}\rho_{n}(t)-\rho_{n}(t)(\mathcal{H}^{\rm eff}_{n})^{\dagger}\right]$ (49)
$+$ $L_{hg,n}^{\rm eff}\rho(t)(L_{hg,n}^{\rm eff})^{\dagger}+L_{eg,n}^{\rm eff}\rho(t)(L_{eg,n}^{\rm eff})^{\dagger}.$
It can be rewritten as a standard master equation
$P_{g}\dot{\rho}_{n}(t)P_{g}=$ $-i\left[H^{\text{eff}}_{n},\rho_{n}(t)\right]+\mathcal{L}[L_{hg,n}^{\rm eff}]\rho_{n}(t)$ (50)
$+\mathcal{L}[L_{eg,n}^{\rm eff}]\rho_{n}(t)$
with a Hermitian effective Hamiltonian
$H^{\rm eff}_{n}=-\Omega^{2}\text{Re}\left[\frac{\tilde{\delta}}{\tilde{\Delta}\tilde{\delta}-g^{2}A_{n}^{2}}\right]|gn\rangle\langle gn|.$ (51)
It is the Hermitian part of the operator
$h_{n}^{\rm eff}=\mathcal{H}^{\rm eff}_{n}+\frac{i}{2}\sum_{k=hg,eg}(L^{\rm eff}_{k,n})^{\dagger}L^{\rm eff}_{k,n}$ (52)
$=$ $\mathcal{H}^{\rm eff}_{n}+\frac{i}{2}V_{+,n}^{\dagger}(\tilde{H}_{n}^{\dagger})^{-1}\sum_{k=hg,eg,he}L_{k,n}^{\dagger}L_{k,n}\tilde{H}_{n}^{-1}V_{+,n}$
$-\frac{i}{2}V_{+,n}^{\dagger}(\tilde{H}_{n}^{\dagger})^{-1}L_{he,n}^{\dagger}L_{he,n}\tilde{H}_{n}^{-1}V_{+,n}$
$=$ $\mathcal{H}^{\rm eff}_{n}+\frac{1}{2}V_{+,n}^{\dagger}(\tilde{H}_{n}^{\dagger})^{-1}\left(\tilde{H}_{n}^{\dagger}-\tilde{H}_{n}\right)\tilde{H}_{n}^{-1}V_{+,n}$
$-\frac{i}{2}V_{+,n}^{\dagger}(\tilde{H}_{n}^{\dagger})^{-1}L_{he,n}^{\dagger}L_{he,n}\tilde{H}_{n}^{-1}V_{+,n}$
$=$ $\mathcal{H}^{\rm eff}_{n}-\frac{1}{2}V_{+,n}^{\dagger}\left[(\tilde{H}_{n}^{\dagger})^{-1}-\tilde{H}_{n}^{-1}\right]V_{+,n}-\frac{i}{2}(L^{\rm eff}_{he,n})^{\dagger}L^{\rm eff}_{he,n}$
$=$ $H^{\rm eff}_{n}-\frac{i}{2}(L^{\rm eff}_{he,n})^{\dagger}L^{\rm eff}_{he,n},$
where $L_{he,n}=\sqrt{\gamma_{he}}|en\rangle\langle hn|$. The non-Hermitian part represents the leakage from $|gn\rangle$ to $|en\rangle$ with the jump operator $L^{\rm eff}_{he,n}=L_{he,n}\tilde{H}_{n}^{-1}V_{+,n}$. It is negligible comparing to the spontaneous decay $\gamma_{eg}$ under the condition $\Delta,\gamma_{hg}\gg\gamma_{he},\Omega$.
For the uppermost ground-state subspace spanned by $\{|gN\rangle,|hN\rangle\}$, the non-Hermitian Hamiltonian reads
$\tilde{H}_{N}=\tilde{\Delta}|hN\rangle\langle hN|.$ (53)
Following the same procedure, the effective Lindblad operator and the effective Hamiltonian turn out to be
$L^{\rm eff}_{hg,N}=$ $L_{hg,N}\tilde{H}_{N}V_{+,N}=\frac{\sqrt{\gamma_{hg}}\Omega}{\tilde{\Delta}}|gN\rangle\langle gN|,$ (54)
$H^{\rm eff}_{N}=$ $-\Omega^{2}{\rm Re}\left(\frac{1}{\tilde{\Delta}}\right)|gN\rangle\langle gN|,$ (55)
respectively. Up to this point, we have established the effective master equation describing the main unidirectional charging mechanism in Fig. 2(a).
The leading-order correction induced by the decay $\gamma_{he}$ is described in Fig. 2(b). The process $|gn\rangle\rightarrow|hn\rangle\rightarrow|en\rangle\rightarrow|gn\rangle$ gives rise to a pure dephasing effect on $|gn\rangle$, which is associated with a higher order term $\mathbf{L}^{(4)}_{{\rm eff},n}$ in Eq. (42)
[…]
×$\times\int_{0}^{t_{1}}dt_{2}\mathcal{V}_{n}(t_{2})\int_{0}^{t_{2}}dt_{3}\mathcal{V}_{n}(t_{3})\mathcal{P}_{g},$
where $\mathcal{J}_{he,n}(t)$ describes the decay $\gamma_{he}$:
$\mathcal{J}_{he,n}(t)=e^{-i\tilde{H}_{n-1}^{*}t}L_{he,n}e^{i\tilde{H}_{n}^{*}t}\otimes e^{i\tilde{H}_{n-1}t}L_{he,n}e^{-i\tilde{H}_{n}t}$ (57)
with $L_{he,n}=\sqrt{\gamma_{he}}|en\rangle\langle hn|$. The dephasing rate turns out to be
$\frac{|\langle gn|J^{\rm eff}_{eg,n}|gn\rangle|^{2}}{|\langle gn|L^{\rm eff}_{hg,n}|gn\rangle|^{2}}=\frac{\gamma_{eg}\gamma_{he}}{\gamma_{hg}}\mathbf{H}^{-1}_{en,en},$ (58)
where $\mathbf{H}^{-1}_{en,en}$ is defined as
$\mathbf{H}^{-1}_{en,en}=\langle en|\otimes\langle en|$ (59)
×$\times\left(-i\tilde{H}^{*}_{n-1}\otimes\mathbb{I}+\mathbb{I}\otimes i\tilde{H}_{n-1}\right)^{-1}|en\rangle\otimes|en\rangle$
using the identity $e^{-i\tilde{H}^{*}t}\otimes e^{i\tilde{H}t}=\exp(-i\tilde{H}^{*}t\otimes\mathbb{I}+\mathbb{I}\otimes i\tilde{H}t)$. Similarly, the process $|gn\rangle\rightarrow|hn\rangle\rightarrow|en\rangle\rightarrow|h(n-1)\rangle\rightarrow|g(n-1)\rangle$ in Fig. 2(b) gives rise to an effective transition from $|gn\rangle$ to $|g(n-1)\rangle$, the relevant jump superoperator is given by
$J^{\rm eff*}_{hg,n}\otimes J^{\rm eff}_{hg,n}=\mathcal{P}_{g}\tilde{L}^{*}_{hg,n-1}(t)\otimes\tilde{L}_{hg,n-1}(t)$ (60)
×$\times\int_{0}^{t}dt_{1}\mathcal{J}_{he,n}(t_{1})\int_{0}^{t_{1}}dt_{2}\mathcal{V}_{n}(t_{2})\int_{0}^{t_{2}}dt_{3}\mathcal{V}_{n}(t_{3})\mathcal{P}_{g}.$
The ratio of the effective decay rate from $|gn\rangle$ to $|g(n-1)\rangle$ in Fig. 2(b) to that from $|gn\rangle$ to $|g(n+1)\rangle$ in Fig. 2(a) is
$\frac{|\langle g(n-1)|J^{\rm eff}_{hg,n}|gn\rangle|^{2}}{|\langle g(n+1)|L^{\rm eff}_{eg,n}|gn\rangle|^{2}}=\frac{\gamma_{hg}\gamma_{he}|\tilde{\delta}|^{2}}{\gamma_{eg}g^{2}A_{n}^{2}}\mathbf{H}^{-1}_{h(n-1),en},$ (61)
where $\mathbf{H}^{-1}_{g(n-1),en}$ is given by
$\mathbf{H}^{-1}_{h(n-1),en}=\langle h(n-1)|\otimes\langle h(n-1)|$ (62)
×$\times\left(-i\tilde{H}^{*}_{n-1}\otimes\mathbb{I}+\mathbb{I}\otimes i\tilde{H}_{n-1}\right)^{-1}|en\rangle\otimes|en\rangle.$
Under the parametric settings that $\gamma_{eg}\approx\gamma_{he}\ll\gamma_{hg}$, it is important to observe that the dephasing rate given by Eq. (58) is negligible in comparison to that in Eq. (47). Similarly, the effective transition rate from $|gn\rangle$ to $|g(n-1)\rangle$ given by Eq. (61) is negligible in comparison to that from $|gn\rangle$ to $|g(n+1)\rangle$ in Eq. (48) under the condition $\gamma_{eg}\approx\gamma_{he}\ll\gamma_{hg}$ and $\Delta\gg\delta$ due to our optimization condition $g^{2}A_{n}^{2}\approx|\tilde{\Delta}\tilde{\delta}|$.

### v2

@@
 Abstract
-In this work, we propose a stable charging scheme mediated by a dissipative three-level system, which allows a unidirectional energy flow from the external power source to a finite-dimensional quantum battery.
-By virtue of the dissipation, the battery avoids the spontaneous discharging induced by the time-reversal symmetry in any unitary charging scheme.
-Irrespective to the initial state, the battery can be eventually stabilized at the maximal-ergotropy state as long as the charger-battery interaction is present.
-We use a Dyson series of Lindbladian superoperator to obtain an effective master equation to describe the battery dynamics and extract the optimization condition for charging efficiency.
-Also it proves that the Fermi’s golden rule can adapt to the non-Hermitian Hamiltonian.
-We justify the outstanding performance of our charging scheme when applied to the finite-size system with a uniform energy ladder operator, the large spin battery, and the truncated harmonic oscillator battery.
+In this work, we propose a stable charging scheme mediated by a three-level system (qutrit), which renders a unidirectional energy flow from an external power source to an $(N+1)$-dimensional quantum battery.
+By virtue of the qutrit dissipation, the battery avoids the spontaneous discharging induced by the time-reversal symmetry of any unitary-charging scheme.
+Irrespective of the initial state, the battery can be eventually stabilized at the maximal-ergotropy state as long as the charger-battery interaction is present.
+We use a Dyson series of Lindbladian superoperator to obtain an effective master equation for the battery, which is found to be equivalent to the high-order Fermi’s golden rule adapted to the non-Hermitian Hamiltonian and spontaneous decay.
+We extract the optimization condition for charging efficiency and justify it in the finite-size battery with uniform energy splitting, the large spin battery, and the truncated harmonic-oscillator battery.
 I Introduction
-As a microscopic unit for energy storage, quantum battery Campaioli et al.
-(2024a) has recently received remarkable attention as both a promising energy supply device for quantum technology Auffèves (2022) and a tool to develop the wide application of quantum thermodynamics Horodecki and Oppenheim (2013); Hovhannisyan et al.
+Quantum battery is a microscopic unit for storing energy Campaioli et al.
+(2024a).
+Recently, it has received remarkable attention as a promising energy supply device for quantum technology Auffèves (2022) and a tool to develop a wide range of applications of quantum thermodynamics Horodecki and Oppenheim (2013); Hovhannisyan et al.
 (2013); Vinjanampathy and Anders (2016); Campaioli et al.
 (2018).
-Since formally introduced by Alicki and Fannes Alicki and Fannes (2013), a lot of efforts have been devoted to studying the role of quantum resources such as quantum correlation and quantum coherence in the charging performance of various quantum batteries Alicki and Fannes (2013); Campaioli et al.
+Since formally introduced by Alicki and Fannes Alicki and Fannes (2013), much effort has been devoted to studying the role of quantum resources, e.g., correlation, coherence, and entanglement, in the charging performance of various quantum batteries Alicki and Fannes (2013); Campaioli et al.
 (2017); Kamin et al.
@@
 (2024).
-It is found that the collective charging schemes could achieve a super-extensive scaling in power, which supports that the quantum correlation provides a dramatic boost to the charging efficiency of the parallel charging schemes Ferraro et al.
+The collective charging schemes could achieve a superextensive scaling in power Ferraro et al.
 (2018); Le et al.
@@
 (2020).
-Quantum speed up in charging can also be found in a single high-dimensional quantum battery via repeated collisions between chargers and battery Seah et al.
+It is demonstrated that the quantum correlation provides a dramatic boost to the charging efficiency of the parallel schemes.
+Repeated collisions between chargers and battery speeds up the charging rate of a finite-dimensional quantum battery Seah et al.
 (2021); Salvia et al.
@@
 (2024).
-In addition, quantum measurements on chargers can give rise to a more efficient charging performance than the measurement-free schemes and do not invoke the initial quantum coherence in both battery and chargers Yan and Jing (2023).
+Quantum measurements on chargers could give rise to a more efficient charging performance than the measurement-free schemes and do not take advantage of the initial coherence in both battery and chargers Yan and Jing (2023).
+In addition, an implementation-independent approach indicates that the charging power is not an entanglement monotone Gyhm and Fischer (2024).
 Numerous prototypes of quantum batteries have been developed in diverse models.
-The one-dimensional spin chain battery has been proposed in Ref.
+A one-dimensional spin chain battery has been proposed in Ref.
 Le et al.
 (2018).
-Then many investigations further extended the spin chain to the spin-ensemble systems Qi and Jing (2021); Huangfu and Jing (2021); Zhao et al.
+Afterwards many investigations extended the spin chain Grazi et al.
+(2025) to the spin-ensemble systems Qi and Jing (2021); Huangfu and Jing (2021); Zhao et al.
 (2021); Arjmandi et al.
@@
 (2022); Ghosh and Sen (De); Yao and Shao (2022).
-For example, by mediating between the charger and battery both consisting of a spin ensemble through a magnon system, we can realize a long-range charging Qi and Jing (2021).
-Dick model battery is another type of extensively studied model Ferraro et al.
+For example, by mediating between the charger and battery both consisting of a spin ensemble through a magnon system, one can realize a long-range and full charging Qi and Jing (2021).
+Dicke-model battery is also under extensive studies Ferraro et al.
 (2018); Andolina et al.
@@
 (2020a, b), which is based on the interaction between an ensemble of noninteracting two-level systems (TLSs) and a common cavity mode.
-The connection and distinction between the spin-charger and cavity-charger schemes were discussed in Ref.
+The similarity and distinction between the spin-charger and cavity-charger schemes were discussed in Ref.
 Huangfu and Jing (2021).
 In addition to the spin systems, the bosonic quantum battery based on an ensemble of harmonic oscillators (HOs) was proposed in Ref.
-Friis and Huber (2018) focusing on the charging precision, i.e., the variance of high-level population distribution.
-Besides the charging efficiency, stabilizing a quantum battery at the charged (high-ergotropy) state is also crucial for the reliable operation over quantum batteries.
-Many conventional schemes have to immediately decouple the charger from the battery of a finite size when the latter reaches the fully charged state.
-If they remain coupled with each other, energy can flow back from the battery to the charger, degrading the charging-efficiency and the energy retention Ferraro et al.
+Friis and Huber (2018), focusing on the charging precision, i.e., the variance of high-level population distribution.
+Besides the charging efficiency, stabilizing the battery system at the highly charged (high-ergotropy) state is also crucial for reliable quantum batteries.
+Many conventional schemes have to immediately decouple the charger from the battery when the latter is fully charged.
+Otherwise, the stored energy can flow back from the battery to the charger, degrading the charging performance and the energy retention Ferraro et al.
 (2018); Zhang et al.
@@
 (2024).
-Extra operations have been designed to address this challenge.
-The Zeno-protection scheme applies sequential projective measurements on the high energy state of a two-level quantum battery Gherardini et al.
+Extra operations have been designed to address discharging.
+Zeno-protection scheme applies sequential projective measurements on the upper state of a two-level quantum battery Gherardini et al.
 (2020).
-The time-dependent Hamiltonian schemes use the stimulated Raman adiabatic passage to maintain a stable charged state Santos et al.
+Selective weak measurements were implemented to protect the battery from discharging by decoherence without extra net recharging Malavazi et al.
+(2025).
+The stimulated Raman adiabatic passage was used to maintain a stable charged state Santos et al.
 (2019); Rojo-Francàs et al.
@@
 Nonlocal and strong chaotic correlation was demonstrated in a Sachdev-Ye-Kitaev quantum battery Rosa et al.
-(2020) to improve the charging stability, to name a few.
+(2020) to improve the charging stability.
 An alternative approach for stable charging is assisted by a dissipative charger connecting the battery and the external power source Farina et al.
+(2019), where both charger and battery are identical HOs or TLSs.
+The energy and ergotropy of the battery can be stabilized with a certain amount determined by the charger-battery coupling strength and the intensity of the external power source.
+However, the battery could not be maintained at its highest energy level in such a setting.
+Introducing a common dissipative environment for battery and charger might induce a unidirectional energy flow and increase the steady energy in the HO battery Ahmadi et al.
 (2024).
-In that model that both charger and battery are identical HOs or TLSs, the energy and ergotropy of the battery can be stabilized at a finite value determined by the charger-battery coupling strength and the intensity of the external power source.
-However, the battery could not be stabilized at its highest energy level in such a setting.
-Introducing a common dissipative environment for the composite system of battery and charger and carefully designing their inner coupling strength and the dissipative rates can increase the steady energy in the battery Ahmadi et al.
-(2024).
+While that scheme was much sensitive to the subtle balance between the charger-battery coupling strength and their dissipative rates.
 In this work, we introduce an efficient and stable charging model that the charger is modeled as a dissipative $\Delta$-type qutrit mediating between the external power source and the battery.
-The charger is coupled to the battery through the energy exchange interaction.
-By deriving an effective master equation for the battery using a perturbation technique for Lindbliadian superoperator, we find that a unidirectional energy flow from the external source to the battery is constructed due to the time-reversal symmetry breaking by the dissipative charger.
-We first consider the battery with a uniform energy ladder operator that is popular in many worksSeah et al.
-(2021); Chen et al.
-(2022); Mitchison et al.
-(2021).
-Our scheme is characterized with universality in enhancing and holding the battery energy for any initial state Sathe and Caravelli (2024) and does not require the charger to be decoupled from the battery after the charging is completed.
-We apply our scheme to more realistic scenarios by modeling quantum battery as a large spin and a truncated quantum harmonic oscillator.
-It is interesting to find that choosing an appropriate charger-battery coupling strength or adjusting it during the charging process can significantly boost the charging efficiency.
+The charger is coupled to the battery through the exchange interaction.
+By deriving an effective master equation about the battery using a perturbation technique for Lindbladian superoperator, we find that a unidirectional energy flow from the external source to the battery is constructed due to the time-reversal symmetry broken by the dissipative charger.
+Our scheme is universal in enhancing and holding the battery energy for any initial state Sathe and Caravelli (2025) and needs not decouple the charger from the battery after the charging is completed.
+We apply it to three battery prototypes with various coupling strengths.
+It is interesting to find that choosing an appropriate charger-battery coupling strength or quenching it during the charging process can significantly boost the charging efficiency.
 The rest of the paper is organized as follows.
@@
 In Sec.
-III, we derive an effective master equation for the battery in a weak-driving regime (see appendix A for details) and obtain the optimization condition for charging efficiency.
-In a broader perspective, we actually develop an efficient technique to extend the Fermi’s Golden rule to adapt to the non-Hermitian Hamiltonian.
+III, we derive an effective master equation for the battery in a weak-driving regime (see appendix A for details) and obtain the optimization condition for charging the uniform-splitting battery.
+In a broader perspective, we develop an efficient technique to extend the Fermi’s golden rule to the situation with non-Hermitian Hamiltonian (battery) and spontaneous decay (charger).
 In Sec.
-IV, we discuss the large spin quantum battery.
+IV, we discuss the large-spin battery.
 It is found that charging efficiency can be significantly boosted by optimizing the effective transition rate for special subspaces.
 In Sec.
-V, we discuss the truncated harmonic oscillator quantum battery.
-It is shown that by adjusting the charger-battery coupling strength according to the instantaneous state of the battery, the charging efficiency can be greatly enhanced.
+V, we discuss the truncated harmonic-oscillator battery.
+It is shown that by adjusting the charger-battery coupling strength due to the instantaneous state of the battery, the charging efficiency can be greatly enhanced.
 The conclusion is provided in Sec.
@@
 In our charging scheme, the dissipative $\Delta$-type qutrit serves as the charger, that is driven by an external power source and transfers energy to the battery.
-Our indirect charging model consists of a quantum battery of a finite size and a dissipative charger that is driven by an external power source.
-As shown in Fig.
-1, the charger is a dissipative $\Delta$-type qutrit.
+Our indirect charging model consists of a quantum battery and a charger that is driven by an external power source.
+As depicted by Fig.
+1, the quantum battery in Secs.
+II and III has $N+1$ uniform energy levels with a bare Hamiltonian ($\hbar=1$) $H_{B}=E_{B}\sum_{n=0}^{N}n|n\rangle\langle n|$.
+The charger is a dissipative $\Delta$-type qutrit.
 An external field directly drives the transition between levels $|g\rangle$ and $|h\rangle$.
-The quantum battery in Secs.
-II and III has $N+1$ uniform energy levels with a bare Hamiltonian ($\hbar=1$) $H_{B}=E\sum_{n=0}^{N}n|n\rangle\langle n|$.
 The charging process is switched on when the quantum battery couples to the qutrit via the transition $|e\rangle\leftrightarrow|h\rangle$.
@@
 In the rotating frame with respect to
-$U=\exp\left\{i\left[H_{B}+\omega_{L}|h\rangle\langle h|+(\omega_{L}-E)|e\rangle\langle e|\right]t\right\},$ (2)
+$U=\exp\left\{i\left[H_{B}+\omega_{L}|h\rangle\langle h|+(\omega_{L}-E_{B})|e\rangle\langle e|\right]t\right\},$ (2)
 the full Hamiltonian reads
-$H^{\prime}=$ $UHU^{\dagger}-iU\frac{\partial}{\partial t}U^{\dagger}=\Delta|h\rangle\langle h|+\delta|e\rangle\langle e|$ (3)
+$H^{\prime}=$ $UHU^{\dagger}-iU\frac{\partial U^{\dagger}}{\partial t}=\Delta|h\rangle\langle h|+\delta|e\rangle\langle e|$ (3)
 $+\Omega\left(|h\rangle\langle g|+|g\rangle\langle h|\right)+g\left(A^{\dagger}|e\rangle\langle h|+A|h\rangle\langle e|\right)$
-with detunings $\Delta=\omega_{hg}-\omega_{L}$ and $\delta=\omega_{eg}+E-\omega_{L}$.
+with detunings $\Delta=\omega_{hg}-\omega_{L}$ and $\delta=\omega_{eg}+E_{B}-\omega_{L}$.
 When the charger is subject to a local dissipative environment, the evolution of the charger-battery system is governed by the master equation Carmichael (1999):
@@
 Figure 2: (a) Main mechanism of our charging scheme.
-The external power source $\Omega$, the charger-battery interaction $gA_{n}$, and the spontaneous decay $\gamma_{eg}$ induces the transitions $|gn\rangle\rightarrow|hn\rangle$, $|hn\rangle\rightarrow|e(n+1)\rangle$, and $|e(n+1)\rangle\rightarrow|g(n+1)\rangle$, respectively.
-They give rise to an effective transition $|gn\rangle\rightarrow|g(n+1)\rangle$ with a rate $\gamma_{\rm eff}$.
-(b) Leading-order transitions induced by the atomic spontaneous emissions from $|h\rangle$ to $|e\rangle$ and from $|e\rangle$ to $|g\rangle$.
+The external power source $\Omega$, the charger-battery interaction $gA_{n}$, and the spontaneous decay $\gamma_{eg}$ induces the transitions $|gn\rangle\leftrightarrow|hn\rangle$, $|hn\rangle\leftrightarrow|e(n+1)\rangle$, and $|e(n+1)\rangle\rightarrow|g(n+1)\rangle$, respectively.
+They give rise to an effective transition $|gn\rangle\rightarrow|g(n+1)\rangle$ with a rate $\gamma_{n,{\rm eff}}$.
+(b) Leading-order discharging transition induced by the atomic spontaneous emissions from $|hn\rangle$ to $|en\rangle$ and from $|h(n-1)\rangle$ to $|g(n-1)\rangle$.
 The decay channel indicated by $\gamma_{he}$ is the main negative factor for charging.
 In our charging scheme, the mediator qutrit is initially prepared in its ground state $|g\rangle$.
-In the weak-driving regime $\Omega\ll\Delta,\ \gamma_{hg}$, one can suppose that the qutrit subject to a dissipative environment remains mostly in its ground state $\rho(t)\approx P_{g}\rho(t)P_{g}$, where $P_{g}\equiv|g\rangle\langle g|$.
-According to the master equation (4), the main mechanism underlying the indirect charging can be illustrated in the subspaces spanned by $\{|gn\rangle,|hn\rangle,|e(n+1)\rangle,|g(n+1)\rangle\}$ for $n<N$.
-As shown in Fig.
-2(a), the system is driven from $|gn\rangle$ to $|hn\rangle$ through the external power source $\Omega$ and from $|hn\rangle$ to $|e(n+1)\rangle$ via the charger-battery interaction with a coupling strength $gA_{n}$.
-Subsequently, the spontaneous decay $\gamma_{eg}$ induces a transition from $|e(n+1)\rangle$ to $|g(n+1)\rangle$.
-These processes constitute an effective unidirectional transition from $|gn\rangle$ to $|g(n+1)\rangle$, which can be quantitatively described by the effective Lindblad operator in Eq.
-(48):
+In the weak-driving regime, i.e., $\Omega<\gamma_{eg}$ and $\Omega\ll\Delta$, the qutrit subject to a dissipative environment can be supposed to remain mostly in its ground state $\rho(t)\approx P_{g}\rho(t)P_{g}$, where $P_{g}\equiv|g\rangle\langle g|$.
+According to the full master equation (4), the main mechanism underlying the indirect charging can be illustrated in the subspaces spanned by $\{|gn\rangle,|hn\rangle,|e(n+1)\rangle,|g(n+1)\rangle\}$ for $n<N$.
+Figure 2(a) describes the leading-order contribution from the driving Hamiltonian $\Omega(|h\rangle\langle g|+|g\rangle\langle h|)$, the coupling Hamiltonian $g(A^{\dagger}|e\rangle\langle h|+A|h\rangle\langle e|)$, and the spontaneous decay $\mathcal{L}[\sqrt{\gamma_{eg}}|g\rangle\langle e|]$.
+In another word, the three-step process $|gn\rangle\rightarrow|hn\rangle\rightarrow|e(n+1)\rangle\rightarrow|g(n+1)\rangle$ and its correction of higher orders by repeating $|hn\rangle\leftrightarrow|e(n+1)\rangle$ constitute the charging mechanism: $|gn\rangle\rightarrow|g(n+1)\rangle$.
+Meanwhile, the leading-order contribution from the driving $\Omega(|h\rangle\langle g|+|g\rangle\langle h|)$, the decay $\mathcal{L}[\sqrt{\gamma_{he}}|e\rangle\langle h|]$, the charger-battery coupling $g(A^{\dagger}|e\rangle\langle h|+A|h\rangle\langle e|)$, and the decay $\mathcal{L}[\sqrt{\gamma_{hg}}|g\rangle\langle h|]$ is described by Fig.
+2(b).
+Specifically, the four-step process $|gn\rangle\rightarrow|hn\rangle\rightarrow|en\rangle\rightarrow|h(n-1)\rangle\rightarrow|g(n-1)\rangle$ and its correction of higher orders by repeating $|en\rangle\leftrightarrow|h(n-1)\rangle$ constitute the discharging mechanism: $|gn\rangle\rightarrow|g(n-1)\rangle$.
+In regard of the perturbation theory, the discharging process is one order higher than the charging process and then charging dominates discharging.
+This nonreciprocity results in a unidirectional energy flow from the external source to the battery.
+Considering Fig. 2(a) for the $n$th subspace, the charging processes can be quantitatively described by the effective Lindblad operator in Eq. (48):
 $L_{eg,n}^{\rm eff}=\frac{\sqrt{\gamma_{eg}}\Omega gA_{n}}{g^{2}A_{n}^{2}-\tilde{\Delta}\tilde{\delta}}|g(n+1)\rangle\langle gn|$ (5)
 with $\tilde{\Delta}\equiv\Delta-i\gamma_{hg}/2-i\gamma_{he}/2$ and $\tilde{\delta}\equiv\delta-i\gamma_{eg}/2$ the complex detunings.
-Due to the large-detuning condition, the decay from $|hn\rangle$ to $|gn\rangle$ contributes to a pure dephasing of the state $|gn\rangle$ in the leading order.
-The relevant effective Lindblad operator given by Eq.
-(47) reads
-$L_{hg,n}^{\rm eff}=\frac{\sqrt{\gamma_{hg}}\Omega\tilde{\delta}}{\tilde{\Delta}\tilde{\delta}-g^{2}A_{n}^{2}}|gn\rangle\langle gn|.$ (6)
-In addition, the energy shift of $|gn\rangle$ induced by the weak driving and the charger-battery interaction can be described by the effective Hamiltonian
+Due to the large-detuning condition, the spontaneous decay from $|hn\rangle$ to $|gn\rangle$ contributes to a pure dephasing of the state $|gn\rangle$ in the leading order.
+The relevant effective Lindblad operator reads
+$L_{hg,n}^{\rm eff}=\frac{\sqrt{\gamma_{hg}}\Omega\tilde{\delta}}{\tilde{\Delta}\tilde{\delta}-g^{2}A_{n}^{2}}|gn\rangle\langle gn|$ (6)
+as given by Eq. (47). In addition, the energy shift of $|gn\rangle$ induced by the weak driving and the charger-battery interaction can be described by the effective Hamiltonian
 $H_{n}^{\rm eff}=-\Omega^{2}{\rm Re}\left(\frac{\tilde{\delta}}{\tilde{\Delta}\tilde{\delta}-g^{2}A_{n}^{2}}\right)|gn\rangle\langle gn|.$ (7)
 The relevant details can be found in appendix A.
-By tracing out the qutrit from Eq. (4), the effective equation of motion of the battery is given by
+By tracing out the qutrit from the full master equation (4), the effective equation of motion for the battery under charging is given by
 $\dot{\rho}_{B}(t)=-i[H_{\rm eff}^{B},\rho_{B}(t)]+\mathcal{L}[L^{B}_{hg,{\rm eff}}]\rho_{B}(t)+\mathcal{L}[L^{B}_{eg,{\rm eff}}]\rho_{B}(t),$ (8)
@@
 $L^{B}_{hg,{\rm eff}}$ $=\sum_{n=0}^{N}\frac{\sqrt{\gamma_{hg}}\Omega\tilde{\delta}}{\tilde{\Delta}\tilde{\delta}-g^{2}A_{n}^{2}}|n\rangle\langle n|=\frac{\sqrt{\gamma_{hg}}\Omega\tilde{\delta}}{\tilde{\Delta}\tilde{\delta}-AA^{\dagger}g^{2}},$ (10a)
-$L^{B}_{eg,{\rm eff}}$ $=\sum_{n=0}^{N}\frac{\sqrt{\gamma_{eg}}\Omega gA_{n}}{g^{2}A_{n}^{2}-\tilde{\Delta}\tilde{\delta}}|n+1\rangle\langle n|=\frac{\sqrt{\gamma_{eg}}\Omega}{A^{\dagger}Ag^{2}-\tilde{\Delta}\tilde{\delta}}A^{\dagger}.$ (10b)
-are the effective Hamiltonian and Lindblad operators in the battery Hilbert space, respectively. When expanding the coefficients of the effective Lindblad operators in terms of series
-$\langle gn|L_{hg,n}^{\rm eff}|gn\rangle$ $=\sqrt{\gamma_{hg}}\sum_{k=0}^{\infty}\left(\frac{g^{2}A_{n}^{2}}{\tilde{\Delta}\tilde{\delta}}\right)^{k}\frac{\Omega}{\tilde{\Delta}}$ (11)
-$=-\sqrt{\gamma_{hg}}\sum_{k=0}^{\infty}\Omega_{\rm eff}^{(2k)},$
-$\langle g(n+1)|L_{eg,n}^{\rm eff}|gn\rangle$ $=-\sqrt{\gamma_{eg}}\sum_{k=0}^{\infty}\left(\frac{g^{2}A_{n}^{2}}{\tilde{\Delta}\tilde{\delta}}\right)^{k}\frac{gA_{n}}{\tilde{\delta}}\frac{\Omega}{\tilde{\Delta}}$
-$=-\sqrt{\gamma_{eg}}\sum_{k=0}^{\infty}\Omega_{\rm eff}^{(2k+1)},$
-it is interesting to find that $\Omega_{\rm eff}^{(l)}$’s are precisely the effective coupling strengths connecting $|i\rangle=|gn\rangle$ and $|f\rangle=|hn\rangle$ or $|e(n+1)\rangle$ up to the first order in $\Omega$ and $l$ order in $gA_{n}$ with respect to the non-Hermitian Hamiltonian in Eq.
-(31):
-$\tilde{H}_{n}$ $=\tilde{H}_{0,n}+\tilde{V}_{n},$ (12)
+$L^{B}_{eg,{\rm eff}}$ $=\sum_{n=0}^{N-1}\frac{\sqrt{\gamma_{eg}}\Omega gA_{n}}{g^{2}A_{n}^{2}-\tilde{\Delta}\tilde{\delta}}|n+1\rangle\langle n|=\frac{\sqrt{\gamma_{eg}}\Omega}{A^{\dagger}Ag^{2}-\tilde{\Delta}\tilde{\delta}}A^{\dagger}.$ (10b)
+are the effective Hamiltonian and Lindblad operators in the whole Hilbert space of battery, respectively.
+Through expanding the coefficients of the effective Lindblad operators in the $n$th subspace:
+$\langle gn|L_{hg,n}^{\rm eff}|gn\rangle$ $=\frac{\sqrt{\gamma_{hg}}\Omega}{\tilde{\Delta}}\sum_{k=0}^{\infty}\left(\frac{g^{2}A_{n}^{2}}{\tilde{\Delta}\tilde{\delta}}\right)^{k}$
+$=\frac{\sqrt{\gamma_{hg}}}{\tilde{\Delta}}\sum_{k=0}^{\infty}\Omega_{\rm eff}^{(2k+1)},$ (11a)
+$\langle g(n+1)|L_{eg,n}^{\rm eff}|gn\rangle$ $=-\frac{\sqrt{\gamma_{eg}}\Omega gA_{n}}{\tilde{\Delta}\tilde{\delta}}\sum_{k=0}^{\infty}\left(\frac{g^{2}A_{n}^{2}}{\tilde{\Delta}\tilde{\delta}}\right)^{k}$
+$=\frac{\sqrt{\gamma_{eg}}}{\tilde{\delta}}\sum_{k=0}^{\infty}\Omega_{\rm eff}^{(2k+2)},$ (11b)
+it is interesting to find that $\Omega_{\rm eff}^{(l)}$’s are precisely the effective coupling strengths linking $|i\rangle=|gn\rangle$ and $|f\rangle=|hn\rangle$ or $|e(n+1)\rangle$ up to the first order in $\Omega$ and (l1)(l-1) order in $gA_{n}$ with respect to the non-Hermitian Hamiltonian $H_{\rm NH}$ in Eq.
+(27).
+In the $n$th subspace as shown in Fig.
+2(a), we have
+$(H_{\rm NH})_{n}$ $=\tilde{H}_{0,n}+\tilde{V}_{n},$ (12)
 $\tilde{H}_{0,n}$ $=\tilde{\Delta}|hn\rangle\langle hn|+\tilde{\delta}|e(n+1)\rangle\langle e(n+1)|,$
-$\tilde{V}_{n}$ $=gA_{n}(|hn\rangle\langle e(n+1)|+|e(n+1)\rangle\langle hn|).$
-Particularly, they can be obtained by the constrained high-order Fermi’s golden rules Kockum et al. (2017) with the form
+$\tilde{V}_{n}$ $=gA_{n}|hn\rangle\langle e(n+1)|+\Omega|hn\rangle\langle gn|+{\rm H.c.}$
+For a standard perturbative derivation, $(H_{\rm NH})_{n}$ in appendix A is partitioned into the excited-subspace Hamiltonian $\tilde{H}_{n}$ and the weak driving Hamiltonian $V_{n}$, given by Eqs.
+(31) and (32), respectively.
+In Eq.
+(12), it is divided into the diagonal component $H_{0,n}$ and the off-diagonal one $\tilde{V}_{n}$.
+Particularly, $\Omega_{\rm eff}^{(l)}$ can be obtained by the constrained Fermi’s golden rules Kockum et al.
+(2017) in the form
 $\Omega^{(l)}_{\rm eff}$ (13)
 $=$ $\sum_{m_{1},m_{2},\cdots,m_{l}}\frac{\tilde{V}_{n,fm_{l-1}}\cdots\tilde{V}_{n,m_{2}m_{1}}\tilde{V}_{n,m_{1}i}}{(\tilde{E}_{i}-\tilde{E}_{m_{1}})(\tilde{E}_{i}-\tilde{E}_{m_{2}})\cdots(\tilde{E}_{i}-\tilde{E}_{m_{l-1}})},$
-where $\tilde{V}_{n,m_{i}m_{j}}=\langle m_{i}|\tilde{V}_{n}|m_{j}\rangle$ with $m_{i}=hn,e(n+1)$ the intermediate states.
-$\tilde{E}_{m_{i}}=\langle m_{i}|\tilde{H}_{0,n}|m_{i}\rangle$ and $\tilde{E}_{i}=\langle gn|\tilde{H}_{0,n}|gn\rangle=0$ are the diagonal elements of $\tilde{H}_{0,n}$.
-For an even $l$, $\Omega_{\rm eff}^{(l)}$ represents the effective coupling strength for connecting $|gn\rangle$ and $|hn\rangle$ through the process $|gn\rangle\rightarrow|hn\rangle\rightarrow|e(n+1)\rangle\cdots\rightarrow|hn\rangle$.
-For an odd $l$, it represents the effective coupling strength for connecting $|gn\rangle$ and $|e(n+1)\rangle$ through the process $|gn\rangle\rightarrow|hn\rangle\rightarrow|e(n+1)\rangle\cdots\rightarrow|e(n+1)\rangle$.
-Similarly, the coefficient of the effective Hamiltonian (9) is the real part of the summation of all round trips about $|gn\rangle$.
+where $\tilde{V}_{n,m_{i}m_{j}}=\langle m_{i}|\tilde{V}_{n}|m_{j}\rangle$ with $m_{i}=hn,e(n+1)$ the excited states, $\tilde{E}_{m_{i}}=\langle m_{i}|\tilde{H}_{0,n}|m_{i}\rangle$, and $\tilde{E}_{i}=\langle gn|\tilde{H}_{0,n}|gn\rangle=0$.
+For an odd $l$, $\Omega_{\rm eff}^{(l)}$ represents the effective coupling strength for connecting $|gn\rangle$ and $|hn\rangle$ through the process $|gn\rangle\rightarrow|hn\rangle\rightarrow|e(n+1)\rangle\cdots\rightarrow|hn\rangle$.
+For an even $l$, it represents the effective coupling strength for connecting $|gn\rangle$ and $|e(n+1)\rangle$ through the process $|gn\rangle\rightarrow|hn\rangle\rightarrow|e(n+1)\rangle\cdots\rightarrow|e(n+1)\rangle$.
+Similarly, the coefficient of the effective Hamiltonian (7) is the real part of the summation of all round trips about $|gn\rangle$
+$\langle gn|H_{n}^{\rm eff}|gn\rangle=$ ${\rm Re}\left[-\frac{\Omega^{2}}{\tilde{\Delta}}\sum_{k=0}^{\infty}\left(\frac{g^{2}A_{n}^{2}}{\tilde{\Delta}\tilde{\delta}}\right)^{k}\right]$ (14)
+$=$ ${\rm Re}\left[-\frac{\Omega}{\tilde{\Delta}}\sum_{k=0}^{\infty}\Omega_{\rm eff}^{(2k+1)}\right].$
+In parallel to appendix A, Eqs.
+(11)-(14) constitute a straightforward and intuitive method to obtain the effective Lindblad operators (10a) and (10b) and the effective Hamiltonian (9) by taking the decay amplitude $\sqrt{\gamma_{hg}}$ and $\sqrt{\gamma_{eg}}$, the Rabi frequency $\Omega$, and the effective splittings $\tilde{\Delta}$ and $\tilde{\delta}$ of all the subspaces into account.
+Alternatively, they give rise to the same effective master equation (8).
 The effective transition rate about $|gn\rangle\rightarrow|g(n+1)\rangle$ in Fig. 2(a) is given by
@@
 $=\frac{\gamma_{eg}\Omega^{2}}{g^{2}A_{n}^{2}+|\tilde{\Delta}\tilde{\delta}|^{2}/(g^{2}A_{n}^{2})-2|\tilde{\Delta}\tilde{\delta}|\cos\phi},$
-where $\phi=\arg(\tilde{\Delta}\tilde{\delta})$.
-With fixed detunings $\tilde{\Delta}$ and $\tilde{\delta}$, one can find that the effective transition rate in the $n$th subspace can be optimized by setting
-$g^{2}_{\rm opt}A^{2}_{n}=|\tilde{\Delta}\tilde{\delta}|.$ (15)
-The decay channel $|h\rangle\rightarrow|e\rangle$ plays a negative role in the charging efficiency through a higher-order process than $|e\rangle\rightarrow|g\rangle$ and $|h\rangle\rightarrow|g\rangle$ (see Appendix A).
-As the leading-order contribution depicted in Fig.
-2(b), the system in the subspace is driven by the external power from $|gn\rangle$ to $|hn\rangle$ then decay from $|hn\rangle$ to $|en\rangle$ and then decay from $|en\rangle$ to $|gn\rangle$ will give rise to an effective dephasing on the state $|gn\rangle$ with a rate given by Eq.
-(58).
-In a biased fluxonium system at about $0.05\leq\Phi_{\rm ex}/\Phi_{0}\leq 0.15$, where $\Phi_{\rm ex}$ and $\Phi_{0}$ are the applied flux and flux quantum, respectively, $\gamma_{he}$ is the same order in magnitude as $\gamma_{eg}$ and is one order of magnitude smaller than $\gamma_{hg}$ Joo et al.
-(2010); Manucharyan et al.
-(2009), i.e.
-$\gamma_{eg}\approx\gamma_{he}\ll\gamma_{hg}$.
-With such a parametric setting, one can confirm that the dephasing rate caused by the process in Fig.
-2(b) is negligible in comparison to that described by Eq.
-(6) for $|gn\rangle\rightarrow|hn\rangle\rightarrow|gn\rangle$ in Fig.
-2(a).
-Similarly, the process that from $|gn\rangle$ to $|hn\rangle$ through the external driving, from $|hn\rangle$ to $|en\rangle$ through the spontaneous decay $\gamma_{he}$, from $|en\rangle$ to $|h(n-1)\rangle$ through the charger-battery interaction, and finally from $|h(n-1)\rangle$ to $|g(n-1)\rangle$ through the spontaneous decay $\gamma_{hg}$ gives rise to an effective transition from $|gn\rangle$ to $|g(n-1)\rangle$ with a rate given by Eq.
-(61).
-It is also negligible comparing to the effective transition rate in Eq.
-(14) under the condition $|\tilde{\Delta}|\gg|\tilde{\delta}|$ and $\gamma_{eg}\approx\gamma_{he}$.
-Consequently, the presence of a decay channel $|h\rangle\rightarrow|e\rangle$ does not significantly alter our unidirectional energy flow.
-Figure 3: Comparison between the dynamics obtained by the full master equation (4) and the effective master equation (8) of (a) battery energy $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ with various $\Omega$.
-The coupling strength $g$ satisfies the optimized condition (15).
-The battery size is set as $N=50$.
-$\gamma_{he}=0$, $\gamma_{hg}=10\gamma_{eg}=0.1E_{B}$, and $\Delta=10\delta=0.1E_{B}$.
-Figure 4: Charging performance in the presence of the decay from $|h\rangle$ to $|e\rangle$ with $\gamma_{he}=\gamma_{eg}$ under various inverse temperature $\beta$ ($k_{B}\equiv 1$).
-(a) Energy stored in the battery $\Delta E(t)$.
-(b) Ergotropy $\mathcal{E}(t)$.
-The driving intensity is fixed as $\Omega=0.05\gamma_{hg}$.
-Other parameters are the same as in Fig.
-3.
-The justification of our charging scheme begins with a simple case with $A_{n}=1$, resulting in uniform effective transition rates across all the subspaces.
+[…]
+[…]
+[…]
+The justification of our charging scheme begins with a battery of a uniform energy ladder configuration, i.e., $A_{n}=1$ that is popular in literature Seah et al.
+(2021); Chen et al.
+(2022); Mitchison et al.
+(2021), resulting in an $n$-independent effective transition rate across all the subspaces.
 Substituting the optimized condition (16) to Eq.
-(14), it is found that the optimized effective transition rate reads
+(15), it is found that the optimized rate reads
 $\gamma^{\rm opt}_{\rm eff}=\frac{\gamma_{eg}\Omega^{2}}{4|\tilde{\Delta}\tilde{\delta}|\sin^{2}(\phi/2)}.$ (17)
@@
 and the ergotropy Allahverdyan et al. (2004), the maximum amount of work that can be extracted from the battery via unitary transformation, which is defined as
-$\mathcal{E}(t)\equiv{\rm Tr}[H_{B}(\rho_{B}(t))]-{\rm Tr}[H_{B}\sigma(t)],$ (18)
-Here $\sigma(t)$ is the passive state that is diagonal in the battery eigenstates basis with non-increasing eigenvalues of $\rho_{B}(t)$:
-$\sigma(t)=\sum_{n=0}^{N}\lambda_{n}(t)|n\rangle\langle n|,\quad\lambda_{n+1}(t)\leq\lambda_{n}(t)$ (19)
-Both energy and ergotropy saturate when the battery state approaches the maximal-ergotropy state $\rho_{B}\rightarrow|N\rangle\langle N|$.
+$\mathcal{E}(t)\equiv{\rm Tr}[H_{B}\rho_{B}(t)]-{\rm Tr}[H_{B}\sigma(t)].$ (19)
+Here $\sigma(t)$ is the passive state that is diagonal in the battery eigenbases with nonincreasing eigenvalues of $\rho_{B}(t)$:
+$\sigma(t)=\sum_{n=0}^{N}\lambda_{n}(t)|n\rangle\langle n|,\quad\lambda_{n+1}(t)\leq\lambda_{n}(t).$ (20)
+Both energy and ergotropy saturate when the battery approaches the maximal-ergotropy state $\rho_{B}\rightarrow|N\rangle\langle N|$.
 By virtue of the last term in Eq.
-(8) interpreted by Eq.
-(10b), our charging scheme can hold on that state without decoupling the charger from the battery after the charging process is completed.
-In Fig.
-3, we compare the dynamics of the battery energy $\Delta E(t)$ and ergotropy $\mathcal{E}(t)$ under the full master equation (4) and the effective master equation (8).
-The battery is initially prepared at its ground state $\rho_{B}(0)=|0\rangle\langle 0|$ with vanishing energy and ergotropy.
+(8) or Eq.
+(10b), our charging scheme can hold that state without decoupling the charger from the battery after the charging is completed.
+The first and second terms in Eq.
+(8) do not contribute to the battery energy variation since both the effective Hamiltonian $H_{\rm eff}^{B}$ in Eq.
+(9) and the effective Lindblad operator $L^{B}_{hg,{\rm eff}}$ in Eq.
+(10a) are diagonal in the energy eigenstates.
+Figure 3: Comparison between the dynamics obtained by the full master equation (4) and the effective master equation (8) of (a) battery energy $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ with various $\Omega$.
+(c) Histogram of the population distribution of the battery eigenlevels $\langle n|\rho_{B}(t)|n\rangle$ at various moments when a certain amount of ergotropy $\mathcal{E}(t)$ has been accumulated in the battery under $\Omega=0.05\gamma_{hg}$.
+The coupling strength $g$ satisfies the optimized condition (16).
+The battery size is set as $N=50$.
+$\gamma_{he}=0$, $\gamma_{hg}=10\gamma_{eg}=0.1E_{B}$, and $\Delta=10\delta=0.1E_{B}$.
+In Figs.
+3(a) and (b), we compare the dynamics of the battery energy $\Delta E(t)$ and ergotropy $\mathcal{E}(t)$ under the full master equation (4) and the effective master equation (8).
+The battery is initially prepared at its ground state $\rho_{B}(0)=|0\rangle\langle 0|$ with zero energy and zero ergotropy.
 It is found that for either energy in Fig.
@@
 3(b), the charging efficiency becomes higher under a stronger driving intensity $\Omega$.
-On the other hand, the deviation between the effective dynamics and the full dynamics increases when $\Omega$ becomes larger and larger.
-It is due to the breakdown of the weak driving assumption.
+In the mean time, the deviation between the effective dynamics and the full dynamics is also enlarged with increasing $\Omega$, which is due to the breakdown of the weak-driving assumption.
+An interesting observation is a sudden slope change during the ergotropy dynamics, roughly around $\mathcal{E}(t)\sim 30E_{B}$ as shown in Fig.
+3(b).
+In contrast, the energy increases smoothly during the charging.
+As a unitary-extractable energy, the ergotropy is the gap between the energy of the battery state $\rho_{B}(t)$ and that of the passive state $\sigma(t)$.
+It is then more sensitive than energy to the variation of population distribution.
+Our numerical simulation shows that once the ergotropy exceeds $\mathcal{E}(t)\approx 30.8E_{B}$, the top energy level becomes the most populated one, indicating a dominant inversion in the population distribution.
+As shown in Fig.
+3(c), at the moments when the battery is charged with $\mathcal{E}(t)=25E_{B}$, $30E_{B}$, and $35E_{B}$, the top-level populations are about $0.005$, $0.05$, and $0.15$, respectively.
+The dramatic raise of the top-level population renders a dramatic change in the eigenvalues of the battery density operator $\rho_{B}(t)$ as well as the population histogram.
+Due to the definition in Eq.
+(20), the passive-state energy will stop rising and turn to descend as the top eigenvalue of $\rho_{B}(t)$ has a dramatic increment.
+The ergotropy therefore experiences a subtle yet sudden change in slope around $\mathcal{E}(t)\sim 30E_{B}$.
+On the other hand, the stored energy continues to rise smoothly under charging as it is just the weighted sum of population distribution.
+Figure 4: Charging performance with decay from $|h\rangle$ to $|e\rangle$, $\gamma_{he}=\gamma_{eg}$, under various inverse temperature $\beta$ ($k_{B}\equiv 1$).
+(a) Stored Energy $\Delta E(t)$.
+(b) Ergotropy $\mathcal{E}(t)$.
+$\Omega=0.05\gamma_{hg}$ and the other parameters are the same as in Fig.
+3.
 A practical charging scheme can transform a high-energy yet zero-ergotropy state, e.g., the thermal state, to be a high-energy and high-ergotropy state.
-In Fig.
-4(a) and 4(b), we demonstrate the charging performance for the initial thermal states with various inverse temperatures $\beta$ by the stored energy $\Delta E(t)$ and ergotropy $\mathcal{E}(t)$, respectively.
+In Figs.
+4(a) and (b), we demonstrate the charging performance for the initial thermal states with various inverse temperatures $\beta$.
 $\beta=\infty$ means the ground state.
-It can be found that both energy and ergotropy still monotonically increase with the charging time and ultimately stabilise with their respective maximum values.
-Note high-ergotropy state is different from high-energy state in population distribution.
-Due to the unidirectional transition from the lower Fock states to the higher ones as indicated in Eq.
-(10b), an initially higher-temperature battery state is more convenient to be transformed to be a higher-ergotropy state than the lower-temperature state.
-The final increment of $\Delta E$ for an initially high-temperature state is significantly smaller than that for $\mathcal{E}(t)$.
-Moveover, we take account of a nonvanishing $\gamma_{he}$ in Fig.
+Both energy and ergotropy still monotonically increase and eventually hold on the maximum value.
+Note high-ergotropy states are different from high-temperature states in population distribution.
+The final increment of $\Delta E$ for an initially high-temperature state is significantly smaller than that of $\mathcal{E}(t)$.
+Due to the unidirectional transition from the lower to the higher states as described by Eq.
+(10b), an initially higher-temperature battery state is more convenient to be transformed to a higher-ergotropy state than the lower-temperature state.
+The two-stage behavior of the ergotropy in slope becomes gradually subtle as the battery starts from a higher temperature state [see the yellow and green lines of Fig.
+4(b)].
+Due to the fact that from the beginning, the top level is significantly populated and then the change in the eigenvalues of $\rho_{B}(t)$ is not as dramatic as that from a zero-temperature state [see the blue lines of Fig.
+4(b)].
+In addition, we take account of a finite $\gamma_{he}$ in Fig.
 4.
 In comparison to Fig.
-3, one can find that the effect of the process shown in Fig.
-2(b) is marginal in the charging efficiency, that is consistent with our analysis.
+3, one can find that the discharging depicted in Fig.
+2(b) has a marginal effect on the charging efficiency, that is consistent with our analysis.
 IV Large spin battery
+To implement our charging scheme in a more physically relevant scenario, we consider a $(2J+1)$-dimensional large-spin battery in this section.
+The battery Hamiltonian reads $H_{B}=E_{B}J_{z}=E_{B}\sum_{m=-J}^{J}m|J,m\rangle\langle J,m|$, where $J$ is an integer or half-integer quantum number of energy.
+The charger and the battery interact with each other through $H_{I}=g(J_{+}|e\rangle\langle h|+J_{-}|h\rangle\langle e|)$, where the creation operator of battery is the collective angular momentum operator
+$J_{+}=\sum_{m=-J}^{J}\sqrt{J(J+1)-m(m+1)}|J,m+1\rangle\langle J,m|,$ (21)
+and $J_{-}=J_{+}^{\dagger}$. The effective transition rate from $|J,m\rangle$ to $|J,m+1\rangle$ can be obtained by substituting $A_{m}=\sqrt{J(J+1)-m(m+1)}$ to Eq. (15). We have
+$\gamma_{m,{\rm eff}}=\left|\frac{\sqrt{\gamma_{eg}}\Omega g\sqrt{J(J+1)-m(m+1)}}{g^{2}[J(J+1)-m(m+1)]-\tilde{\Delta}\tilde{\delta}}\right|^{2}.$ (22)
+In contrast to the uniform model with $A_{n}=1$, the effective transition rates in the large spin model vary across different subspaces. The same is true for the optimization condition:
+[…]
 Figure 5: Effective transition rate $\gamma_{m,{\rm eff}}$ (in unit of $E_{B}$) from $|J,m\rangle$ to $|J,m+1\rangle$ in the parameter space of $g$ (in unit of $E_{B}$) and $m$.
@@
 3.
-Figure 6: Full and effective dynamics of (a) battery energy $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ under the optimization conditions of $g=g_{m=-J,{\rm opt}}$ (solid lines and dashed lines) and $g=g_{m=0,{\rm opt}}$ (dotted lines and dot-dashed line), as defined in Eq.
+Figure 6: Full and effective dynamics of (a) battery energy $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ under the optimization conditions of $g=g_{m=-J,{\rm opt}}$ (solid lines and dashed lines) and $g=g_{m=0,{\rm opt}}$ (dotted lines and dot-dashed lines), as defined in Eq.
 (23).
-The quantum number of energy is set as $J=25$ and the decay rate is $\gamma_{he}=\gamma_{eg}$.
-$\Omega=0.1\gamma_{hg}$.
+$J=25$, $\gamma_{he}=\gamma_{eg}$, and $\Omega=0.1\gamma_{hg}$.
 The other parameters are the same as in Fig.
 3.
-To implement our charging scheme in a more physically relevant scenario, we consider an $N=2J+1$ dimensional large-spin battery in this section.
-The battery Hamiltonian reads $H_{B}=E_{B}J_{z}=E_{B}\sum_{m=-J}^{J}m|J,m\rangle\langle J,m|$, where $J$ is an integer or half-integer quantum number of energy.
-The charger and the battery interact with each other through
-$H_{I}=g(J_{+}|e\rangle\langle h|+J_{-}|h\rangle\langle e|),$ (20)
-where the creation operator of battery is the collective angular momentum operator
-$J_{+}=\sum_{m=-J}^{J}\sqrt{J(J+1)-m(m+1)}|J,m+1\rangle\langle J,m|,$ (21)
-and $J_{-}=J_{+}^{\dagger}$. The effective transition rate from $|J,m\rangle$ to $|J,m+1\rangle$ can be obtained by substituting $A_{m}=\sqrt{J(J+1)-m(m+1)}$ to Eq. (14). We have
-$\gamma_{m,{\rm eff}}=\left|\frac{\sqrt{\gamma_{eg}}\Omega g\sqrt{J(J+1)-m(m+1)}}{g^{2}[J(J+1)-m(m+1)]-\tilde{\Delta}\tilde{\delta}}\right|^{2}.$ (22)
-In contrast to the uniform model with $A_{n}=1$, the effective transition rates in the large spin model vary across different subspaces, the same is true for the optimization condition:
-$g_{m,{\rm opt}}^{2}[J(J+1)-m(m+1)]=|\tilde{\Delta}\tilde{\delta}|.$ (23)
-Figure 7: Full and effective dynamics of (a) energy stored in the battery $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ under initial thermal states with different inverse temperature $\beta$.
+In Fig.
+5, we present the dependence of effective transition rate $\gamma_{m,{\rm eff}}$ on the coupling strength $g$ between the qutrit charger and the large-spin battery and the quantum number of angular momentum $m$.
+The optimization condition in Eq.
+(23) is indicated by the black dashed line.
+It is found that by setting $m=0$ in Eq.
+(23), the relevant $\gamma_{m,{\rm eff}}$ closely mimics the optimal conditions across a substantial number of subspaces as shown by the red dash-dotted line.
+In practice, it is an $m$-independent setting for enhancing the charging efficiency.
+In Figs.
+6(a) and 6(b), we compare the charging efficiency under the optimized condition of $g=g_{m=-J,{\rm opt}}$ with that under $g=g_{m=0,{\rm opt}}$ when the battery is initialized as the ground state $\rho_{B}(0)=|J,-J\rangle\langle J,-J|$.
+Under the optimized condition $g=g_{0,{\rm opt}}$, the ergotropy remains almost vanishing during the beginning stage of the charging, while the energy grows with a small rate.
+It is because the effective transition rate from $|J,-J\rangle$ to $|J,-J+1\rangle$ is much smaller than the optimized one from $|J,0\rangle$ to $|J,1\rangle$, i.e., the battery is almost in a passive state $\rho_{B}(t)\approx\sigma(t)$ on the beginning stage.
+As the charging proceeds, the ergotropy becomes nonvanishing when the population inversion around $|J,-J\rangle$ is realized.
+The optimized condition $g=g_{-J,{\rm opt}}$ exhibits an advantage in charging rate on the beginning stage over $g=g_{0,{\rm opt}}$ due to the initial population distribution, i.e., the population on $|J,-J\rangle$ can be efficiently moved to the high level $|J,-J+1\rangle$.
+After a sufficient amount of energy accumulated in the battery, the condition $g=g_{0,{\rm opt}}$ quickly exceeds $g=g_{-J,{\rm opt}}$ since a dominant number of subspaces are efficient in charging (see Fig.
+5).
+Nevertheless, the effective dynamics under $g=g_{0,{\rm opt}}$ exhibits a moderate deviation from the full one and it matches perfectly with the full dynamics under $g=g_{-J,{\rm opt}}$.
+It is due to the tradeoff between the charging efficiency and the validity of the effective master equation under weak driving.
+A higher charging efficiency means the excited levels $|h\rangle$ and $|e\rangle$ of the charger are more populated as indicated by Fig.
+2(a), that disrupts the approximation $\rho(t)\approx P_{g}\rho(t)P_{g}$.
+On the contrary, only the transition $|J,-J\rangle\rightarrow|J,-J+1\rangle$ is efficient under $g=g_{-J,{\rm opt}}$.
+The slow charging then supports the agreement between the effective and full master equations.
+Figure 7: Full and effective dynamics of (a) energy stored in the battery $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ under initial thermal states with various inverse temperatures $\beta$.
 The other parameters are the same as in Fig.
 6.
-In Fig.
-5, we present the dependence of effective transition rate $\gamma_{m,{\rm eff}}$ on the coupling strength between the qutrit charger and the large-spin battery $g$ and the quantum number of angular momentum $m$ to confirm the optimization condition in Eq.
-(23), as indicated by the black dashed line.
-It is noted that by setting $m=0$ in Eq.
-(23), the relevant $\gamma_{m,{\rm eff}}$ closely approximates the optimal conditions across a substantial number of subspaces as shown by the red dash-dotted line.
-In practice, it is a $m$-independent setting for enhancing the charging efficiency.
-In Fig.
-6(a) and 6(b), we compare the charging efficiency under the optimized condition of $g=g_{m=-J,{\rm opt}}$ with that under $g=g_{m=0,{\rm opt}}$ when the battery is initialised as the ground state $\rho_{B}(0)=|J,-J\rangle\langle J,-J|$.
-It is found that while the optimized condition $g=g_{-J,{\rm opt}}$ exhibits an advantage at the beginning of charging due to the initial population distribution, the charging rate under $g=g_{0,{\rm opt}}$ dramatically increases as the charging proceeds.
-The latter quickly surpasses the former.
-This behavior is consistent with our effective master equation.
-Nevertheless, the effective dynamics under $g=g_{0,{\rm opt}}$ exhibits a moderate deviation from the full one, and it perfectly matches with the full dynamics under $g=g_{-J,{\rm opt}}$.
-Now we check the case that the large-spin battery is initially prepared at the thermal states with various inverse temperature $\beta$ by the dynamics of stored energy $\Delta E(t)$ and ergotropy $\mathcal{E}(t)$ in Fig.
-7(a) and 7(b), respectively.
-The optimization condition is chosen as $g=g_{0,{\rm opt}}$.
-It is found that the ergotropy can be maintained about $\mathcal{E}_{\rm steady}>0.995\mathcal{E}_{\rm max}$, irrespective to the temperature $\beta$ of the initial state, where $\mathcal{E}_{\rm max}=2JE_{B}$.
-Moreover, it is interesting to find that a higher initial temperature of the battery results in a higher growth rate of the ergotropy.
+Now we check the large-spin battery prepared as finite-temperature states.
+The dynamics of stored energy $\Delta E(t)$ and ergotropy $\mathcal{E}(t)$ are plotted in Figs.
+7(a) and (b), respectively.
+The optimization condition for charging is chosen as $g=g_{0,{\rm opt}}$.
+The steady ergotropy is held about $\mathcal{E}_{\rm steady}>0.995\mathcal{E}_{\rm max}$ with $\mathcal{E}_{\rm max}=2JE_{B}$, irrespective of $\beta$.
+Moreover, it is interesting to find again that a higher initial temperature of the battery results in a higher growth rate of the ergotropy.
 V Harmonic oscillator battery
@@
 Subsequently, it can be optimized under the condition
-$g_{n,{\rm opt}}^{2}(n+1)=|\tilde{\Delta}\tilde{\delta}|.$ (25)
-Figure 9: Full and effective dynamics of (a) battery energy $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ with an invariant coupling strength $g=g_{0,{\rm opt}}$ in Eq.
-(25) during the charging process (black solid lines and blue dashed lines) and with a coupling strength quenched at $\gamma_{eg}\tau$ (green and yellow lines) according to Eq.
+[…]
+The effective transition rate in the space of the coupling strength $g$ and the subspace index $n$ is plotted in Fig.
+8.
+The optimization condition is indicated by the black dashed line.
+Distinct from the large-spin battery, it is hardly to find a nearly subspace-independent coupling strength $g_{\rm opt}$ to optimize the effective transition rate for the HO battery, since $g_{n,{\rm opt}}$ decreases monotonically with $n$.
+Nevertheless, the charging efficiency can be enhanced by quenching the coupling strength during the charging process.
+By our scheme, one can initially choose $g=g_{0,{\rm opt}}$ since the ground level is always the most occupied one for a thermal state.
+Then after a duration $\tau$, the coupling strength can be quenched as
+$g=g_{\bar{n}(\tau),{\rm opt}}=\sqrt{\frac{|\tilde{\Delta}\tilde{\delta}|}{\bar{n}(\tau)+1}},$ (26)
+where $\bar{n}(\tau)={\rm Tr}[\rho_{B}(\tau)H_{B}]/E_{B}$ is the instantaneous average excitation of the truncated HO battery.
+Figure 9: (a) Saturation time $\gamma_{eg}T$ of ergotropy versus the quench time $\gamma_{eg}\tau$.
+(b) Effective dynamics of the ergotropy $\mathcal{E}(t)$ under various settings of quenching.
+The coupling strength is initially $g=g_{0,{\rm opt}}$ in Eq.
+(25) and then quenched according to Eq.
 (26).
@@
 3.
-The effective transition rate in the parameter space of charger-battery coupling strength $g$ and $n$ is plotted in Fig.
-8.
-The optimization condition is indicated by the black dashed line.
-Distinct from the large-spin battery, it is hardly to find a subspace-independent coupling strength $g_{\rm opt}$ to optimize the effective transition rates for the harmonic-oscillator battery, since $g_{n,{\rm opt}}$ decreases with $n$ in a monotonic way.
-Nevertheless, the charging efficiency can be enhanced by adjusting the coupling strength during the charging process.
-By our scheme, one can initially optimize the effective transition rate from $|0\rangle$ to $|1\rangle$ by choosing $g=g_{0,{\rm opt}}$ in Eq.
-(25) due to the fact that the ground state is always the most populated state for a thermal state.
-Then after an interval of $\tau$, one can modify the coupling strength to be
-$g=g_{\bar{n}(\tau),{\rm opt}}=\sqrt{\frac{|\tilde{\Delta}\tilde{\delta}|}{\bar{n}(\tau)+1}},$ (26)
-where $\bar{n}(\tau)={\rm Tr}[\rho_{B}(\tau)H_{B}]/E_{B}$ is the instantaneous average excitation of the truncated HO battery.
-We compare the charging performance with a uniform $g$ during the whole charging evolution and that with a quenched $g$ by the dynamics of $\Delta E(t)$ and $\mathcal{E}(t)$ in Figs.
-9(a) and 9(b), respectively.
-The yellow and green lines indicate the quenched moment at $\gamma_{eg}\tau=500$ and $\gamma_{eg}\tau=1000$, respectively, in which the charger-battery coupling strength is only modified at the specified moment and calculated according to Eq.
-(26).
-The battery is initially prepared in the ground state $\rho_{B}(0)=|0\rangle\langle 0|$.
-It is clearly that the strategy of a uniform $g$ is the lowest in charging efficiency and both energy and ergotropy have an abrupt and significant enhancement when $g$ can be re-optimized.
-The numerical simulation indicates that the full dynamics is perfectly matched with the effective dynamics and the charging efficiency is slightly overestimated after the adjustment of the coupling strength $g$.
-Figure 10: Full and effective dynamics of (a) energy stored in the battery $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ under initial thermal states with various $\beta$.
+In Fig.
+9(a), we demonstrate the saturation time $\gamma_{eg}T$ for charging the truncated HO battery to the top energy level as a function of the quench time $\gamma_{eg}\tau$.
+It is found that $\gamma_{eg}T$ can have a comparably small and steady value in the range of $500\leq\gamma_{eg}\tau\leq 1000$.
+In addition, the saturation time of charging can be further reduced if one can apply multiple quenches within that range of quench time.
+As shown in Fig.
+9(b), the quench effect is quickly saturated.
+Figure 10: Full and effective dynamics of (a) battery energy $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ with a fixed coupling strength $g=g_{0,{\rm opt}}$ during the charging process (black solid lines and blue dashed lines) and with a coupling strength quenched different moments (green and yellow lines).
+The other parameters are the same as in Fig.
+9.
+Figure 11: Full and effective dynamics of (a) stored energy in the battery $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ under initial thermal states with various $\beta$.
 The coupling strength is only modified at $\gamma_{eg}\tau=500$ according to Eq.
@@
 The other parameters are the same as in Fig.
-9.
-The dynamics of the stored energy $\Delta E(t)$ and ergotropy $\mathcal{E}(t)$ for charging the HO battery initially in the thermal states with a finite temperature is demonstrated in Figs.
-10(a) and 10(b), respectively.
+10.
+We present the charging performance with a fixed $g$ and that with a quenched $g$ by the dynamics of $\Delta E(t)$ in Fig.
+10(a) and $\mathcal{E}(t)$ in Fig.
+10(b).
+The yellow and green lines indicate the quenched moment at $\gamma_{eg}\tau=500$ and $\gamma_{eg}\tau=1000$, respectively.
+The battery is initially prepared in the ground state $\rho_{B}(0)=|0\rangle\langle 0|$.
+Apparently, the dynamics of both energy and ergotropy have an abrupt and dramatic acceleration when $g$ can be re-optimized.
+The numerical simulation indicates that the full dynamics perfectly matches with the effective dynamics and the charging efficiency is slightly overestimated after the adjustment of the coupling strength $g$.
+The dynamics of the stored energy $\Delta E(t)$ and ergotropy $\mathcal{E}(t)$ for charging the HO battery initially in the finite-temperature states is demonstrated in Figs.
+11(a) and (b), respectively.
 The results are similar to the large-spin battery presented in Fig.
@@
 A higher initial temperature or energy results in a higher growth rate of the ergotropy.
-In addition, the steady value of the ergotropy is insensitive to the initial inverse temperature $\beta$.
+Again, the steady value of the ergotropy is insensitive to the initial inverse temperature $\beta$.
 In this case, we have $\mathcal{E}_{\rm steady}>0.994\mathcal{E}_{\rm max}$ with $\mathcal{E}_{\rm max}=NE_{B}$.
 VI Conclusion
-In summary, we have proposed a stable charging scheme assisted by a dissipative $\Delta$-type qutrit, which relies on the nonreciprocal interaction between the external power source and the battery and the ensued unidirectional energy flow.
-Three types of finite dimensional quantum batteries, including the uniform coupling strength model, the large-spin battery, and the truncated harmonic oscillator battery, have confirmed that the battery energy and ergotropy can be maintained at their respective maximum values and the battery need not to be decoupled from the charger after the charging process is completed.
-By deriving the effective master equation of the battery, we can obtain the optimization condition of the charging efficiency.
-For the large spin battery, the charging efficiency can be significantly enhanced by choosing the charger-battery coupling strength according to the optimization condition in the subspace with $m=0$.
-For the truncated harmonic oscillator battery, the enhancement can be achieved by adjusting the coupling strength according to the instantaneous state of the battery during charging.
+In summary, we proposed a stable charging scheme assisted by a dissipative qutrit yielding an unidirectional energy flow, which is based on the effective nonreciprocal interaction between the external power source and the battery.
+Three prototypes of finite dimensional quantum batteries, including the uniform coupling-strength model, the large-spin battery, and the truncated harmonic oscillator battery, have confirmed that the battery energy and ergotropy can hold on the maximum values and the battery need not to be decoupled from the charger after the charging process is completed.
+By deriving the effective master equation for the battery, that could be alternatively obtained by the generalized Fermi’s golden rule adapted to non-Hermitian Hamiltonian and spontaneous decay, one can extract the optimized condition of the charging rate.
+For the large-spin battery, the charging efficiency can be significantly enhanced by choosing the charger-battery coupling strength according to the optimization condition in the subspace indexed with $m=0$.
+For the truncated harmonic oscillator battery, the enhancement can be achieved by quenching the coupling strength according to the instantaneous state of the battery.
 In application, our work provides an efficient initial-state-independent charging scheme and avoids the discharging induced by the time-reversal symmetry.
-In theory, it extends the high-order Fermi’s Golden rule to the regime of non-Hermitian Hamiltonian.
-Appendix A The effective master equation
+Acknowledgments
+We acknowledge grant support from the Science and Technology Program of Zhejiang Province (No. 2025C01028).
+Appendix A Effective master equation
 This appendix contributes to deriving the effective Hamiltonian and Lindblad operators for the quantum battery $\rho_{B}(t)$ using a technique extending the James’ method James and Jerke (2007); Gamel and James (2010); Shao et al.
 (2017) from Hamiltonian to Lindbladian.
+The fundamental idea is still based on the large detuning condition and the relevant rotating-wave approximation.
 By defining a non-Hermitian Hamiltonian
@@
 where $L_{hg}$, $L_{eg}$, and $L_{he}$ are $\sqrt{\gamma_{hg}}|g\rangle\langle h|$, $\sqrt{\gamma_{eg}}|g\rangle\langle e|$, and $\sqrt{\gamma_{he}}|h\rangle\langle e|$, respectively, and $\tilde{\Delta}=\Delta-i\gamma_{hg}/2-i\gamma_{he}/2$ and $\tilde{\delta}=\delta-i\gamma_{eg}/2$ are complex detunings, the full master equation (4) can be rewritten as
-$\dot{\rho}(t)=$ $-i[(\tilde{H}+V)\rho(t)-\rho(t)(\tilde{H}^{\dagger}+V)]$ (30)
+$\dot{\rho}(t)=$ $-i\left[(\tilde{H}+V)\rho(t)-\rho(t)(\tilde{H}^{\dagger}+V)\right]$ (30)
 $+L_{hg}\rho(t)L_{hg}^{\dagger}+L_{eg}\rho(t)L_{eg}^{\dagger}+L_{he}\rho(t)L_{he}^{\dagger}.$
 In the weak-driving regime Reiter and Sørensen (2012), the subspace spanned by $\{|gn\rangle,|hn\rangle,|e(n+1)\rangle,|g(n+1)\rangle\}$ with $n<N$ in Fig.
-2(a) can be divided into the rapidly-decaying exited subspace spanned by $\{|hn\rangle,|e(n+1)\rangle\}$ and the ground subspace spanned by $\{|gn\rangle,|g(n+1)\rangle\}$.
+2(a) can be divided into the rapidly decaying exited subspace spanned by $\{|hn\rangle,|e(n+1)\rangle\}$ and the ground subspace spanned by $\{|gn\rangle,|g(n+1)\rangle\}$.
 With respect to Fig.
@@
 $+gA_{n}\left[|hn\rangle\langle e(n+1)|+|e(n+1)\rangle\langle hn|\right],$
-involving only with the excited subspaces. The weak external driving $V$ leads to the excitation process from the ground subspaces to the excited subspaces with a Hermitian Hamiltonian
-$V_{n}=V_{+,n}+V^{\dagger}_{+,n}=\Omega|hn\rangle\langle gn|+{\rm H.c.}.$ (32)
+involving only with the excited subspaces. The weak external driving $V$ connects the ground subspaces to the excited subspaces with a Hermitian Hamiltonian
+$V_{n}=V_{+,n}+V^{\dagger}_{+,n}=\Omega|hn\rangle\langle gn|+{\rm H.c.}$ (32)
 The Lindblad operators induced by the qutrit decay read
@@
 To employ the perturbation theory, it is convenient to transform to the rotating frame with respect to $\tilde{U}(t)=e^{i\tilde{H}t}$. The master equation (30) then becomes
-$\dot{\tilde{\rho}}(t)=$ $-i[\tilde{V}(t)\tilde{\rho}(t)-\tilde{\rho}(t)\tilde{V}^{\dagger}(t)]+\tilde{L}_{hg}(t)\tilde{\rho}(t)\tilde{L}^{\dagger}_{hg}(t)$ (34)
+$\dot{\tilde{\rho}}(t)=$ $-i\left[\tilde{V}(t)\tilde{\rho}(t)-\tilde{\rho}(t)\tilde{V}^{\dagger}(t)\right]+\tilde{L}_{hg}(t)\tilde{\rho}(t)\tilde{L}^{\dagger}_{hg}(t)$ (34)
 $+\tilde{L}_{eg}(t)\tilde{\rho}(t)\tilde{L}^{\dagger}_{eg}(t)+\tilde{L}_{he}(t)\tilde{\rho}(t)\tilde{L}^{\dagger}_{he}(t),$
 where $\tilde{\rho}(t)=\tilde{U}(t)\rho(t)\tilde{U}^{\dagger}(t)$ is the density matrix in the rotating frame.
-The other operators are transformed accordingly with $\tilde{\mathcal{O}}(t)=\tilde{U}(t)\mathcal{O}\tilde{U}^{-1}(t)$, $\mathcal{O}=V$, $L_{hg}$, $L_{eg}$, and $L_{he}$.
-With respect to the main process in Fig. 2(a), the transformed master equation (34) in the subspace can be written as
+The other operators are transformed accordingly to be $\tilde{\mathcal{O}}(t)=\tilde{U}(t)\mathcal{O}\tilde{U}^{-1}(t)$, $\mathcal{O}=V$, $L_{hg}$, $L_{eg}$, and $L_{he}$.
+With respect to the main transitions in Fig. 2(a), the transformed master equation (34) in the $n$th subspace can be written as
 $\dot{\tilde{\rho}}_{n}(t)=-i\left[\tilde{V}_{n}(t)\tilde{\rho}_{n}(t)-\tilde{\rho}_{n}(t)\tilde{V}_{n}^{\dagger}\right]$ (35)
-$+\tilde{L}_{hg,n}(t)\tilde{\rho}_{n}(t)\tilde{L}^{\dagger}_{hg,n}(t)$
-$+\tilde{L}_{eg,n}(t)\tilde{\rho}_{n}(t)\tilde{L}^{\dagger}_{eg,n}(t),$
+$+$ $\tilde{L}_{hg,n}(t)\tilde{\rho}_{n}(t)\tilde{L}^{\dagger}_{hg,n}(t)+\tilde{L}_{eg,n}(t)\tilde{\rho}_{n}(t)\tilde{L}^{\dagger}_{eg,n}(t),$
 where
-$\tilde{V}_{n}(t)$ $=\Omega(e^{i\tilde{H}_{n}t}|hn\rangle\langle gn|+|gn\rangle\langle hn|e^{-i\tilde{H}_{n}t}),$ (36)
+$\tilde{V}_{n}(t)$ $=\Omega\left(e^{i\tilde{H}_{n}t}|hn\rangle\langle gn|+|gn\rangle\langle hn|e^{-i\tilde{H}_{n}t}\right),$ (36)
 $\tilde{L}_{hg,n}(t)$ $=\sqrt{\gamma_{hg}}|gn\rangle\langle hn|e^{-i\tilde{H}_{n}t},$ (37)
@@
 The last term about $\tilde{L}_{he}(t)$ in Eq. (34) that appears only as the leading-order contribution depicted in Fig. 2(b) will be analyzed later.
-Then we vectorize the density matrix of each subspace to be a $4^{2}\times 1$ vector denoted by $|\tilde{\rho}_{n}(t)\rangle\rangle$. The equation of motion in this representation is given by
+Then we vectorize the density matrix of each subspace to be a $4^{2}\times 1$ vector denoted by $|\tilde{\rho}_{n}(t)\rangle\rangle$. The equation of motion in the Hilbert-Schmidt space is given by
 $\partial_{t}|\tilde{\rho}_{n}(t)\rangle\rangle=\mathbf{L}_{n}(t)|\tilde{\rho}_{n}(t)\rangle\rangle.$ (39)
-Her, $\mathbf{L}(t)$ is the $4^{2}\times 4^{2}$ superoperator in the form of
+Here $\mathbf{L}(t)$ is the $4^{2}\times 4^{2}$ superoperator in the form of
 $\mathbf{L}_{n}(t)=\mathcal{V}_{n}(t)+\mathcal{J}_{n}(t),$ (40)
-$\mathcal{V}_{n}(t)=-i(\mathbb{I}\otimes\tilde{V}_{n}(t)-\tilde{V}_{n}^{*}(t)\otimes\mathbb{I}),$
+$\mathcal{V}_{n}(t)=-i\left[\mathbb{I}\otimes\tilde{V}_{n}(t)-\tilde{V}_{n}^{*}(t)\otimes\mathbb{I}\right],$
 $\mathcal{J}_{n}(t)=\tilde{L}^{*}_{hg,n}(t)\otimes\tilde{L}_{hg,n}(t)+\tilde{L}^{*}_{eg,n}(t)\otimes\tilde{L}_{eg,n}(t),$
@@
 (2024b).
-$\mathcal{V}(t)$ and $\mathcal{J}(t)$ represent the non-Hermitian evolution part and the quantum jump part, respectively.
+$\mathcal{V}(t)$ and $\mathcal{J}(t)$ represent the non-unitary evolution part and the quantum jump part, respectively.
 The formal solution of Eq.
@@
 $|\tilde{\rho}_{n}(t)\rangle\rangle=|\tilde{\rho}_{n}(0)\rangle\rangle+\int_{0}^{t}dt_{1}\mathbf{L}_{n}(t_{1})|\tilde{\rho}_{n}(t_{1})\rangle\rangle.$ (41)
-In the large detuning condition, i.e., $\Delta+\delta\gg\Omega$, the transformation operator $U(t)=e^{i\tilde{H}_{n}(t)}\propto\exp[i(\tilde{\Delta}+\tilde{\delta})t/2]$ is highly oscillating with time.
+Under the large detuning condition, i.e., $\Delta+\delta\gg\Omega$, the transformation operator $\tilde{U}_{n}(t)=e^{i\tilde{H}_{n}t}\propto\exp[i(\tilde{\Delta}+\tilde{\delta})t/2]$ is highly oscillating with time.
 By iteratively substituting Eq.
@@
 (39), neglecting the fast oscillating terms $\mathcal{V}(t)|\tilde{\rho}_{n}(0)\rangle\rangle$, and taking the Markovian approximation Shao et al.
-(2017), the equation of motion can written into a Dyson series:
+(2017), the equation of motion can written into a perturbative Dyson series:
 $\partial_{t}|\tilde{\rho}_{n}(t)\rangle\rangle=$ $\left[\mathbf{L}_{n}(t)\int_{0}^{t}dt_{1}\mathbf{L}_{n}(t_{1})+\mathbf{L}_{n}(t)\int_{0}^{t}dt_{1}\mathbf{L}_{n}(t_{1})\int_{0}^{t_{1}}dt_{2}\mathbf{L}_{n}(t_{2})+\cdots\right]|\tilde{\rho}_{n}(t)\rangle\rangle.$ (42)
-$=$ $\left[\mathbf{L}^{(2)}_{{\rm eff},n}(t)+\mathbf{L}^{(3)}_{{\rm eff},n}(t)+\cdots\right]|\tilde{\rho}_{n}(t)\rangle\rangle.$
+$=$ $\left[\mathbf{L}^{(2)}_{{\rm eff},n}(t)+\mathbf{L}^{(3)}_{{\rm eff},n}(t)+\mathbf{L}^{(4)}_{{\rm eff},n}+\cdots\right]|\tilde{\rho}_{n}(t)\rangle\rangle.$
 It is assumed that the qutrit remains mostly in its ground state during the charging process $|\tilde{\rho}(t)\rangle\rangle\approx\mathcal{P}_{g}|\tilde{\rho}(t)\rangle\rangle$, where $\mathcal{P}_{g}=P_{g}\otimes P_{g}$.
@@
 $=\mathcal{P}_{g}\bigg{[}\mathcal{V}_{n}(t)\int_{0}^{t}dt_{1}\mathcal{V}_{n}(t_{1})$
-$+\mathcal{J}_{n}(t)\int_{0}^{t}dt_{1}\mathcal{V}_{n}(t_{1})\int_{0}^{t_{1}}dt_{2}\mathcal{V}_{n}(t_{2})\bigg{]}\mathcal{P}_{g}|\tilde{\rho}_{n}(t)\rangle\rangle.$
-up to the second order in driving intensity $\Omega$. By substituting Eq. (36) into the first term of the Eq. (43), we can obtain the effective superoperator for a non-Hermitian evolution:
-[…]
-$=$ $-\left[P_{g}\otimes V_{+,n}^{\dagger}(i\tilde{H}_{n})^{-1}V_{+,n}+V_{+,n}^{\dagger}(-i\tilde{H}_{n}^{*})^{-1}V_{+,n}\otimes P_{g}\right],$
-[…]
+$+\mathcal{J}_{n}(t)\int_{0}^{t}dt_{1}\mathcal{V}_{n}(t_{1})\int_{0}^{t_{1}}dt_{2}\mathcal{V}_{n}(t_{2})\bigg{]}\mathcal{P}_{g}|\tilde{\rho}_{n}(t)\rangle\rangle$
+up to the second order in the driving intensity $\Omega$. By substituting Eq. (36) into the first term of the Eq. (43), we can obtain the effective superoperator for a nonunitary evolution:
+$\mathcal{P}_{g}\mathcal{V}_{n}(t)\int_{0}^{t}dt_{1}\mathcal{V}_{n}(t_{1})\mathcal{P}_{g}$ (44)
+$=$ $-P_{g}\otimes V_{+,n}^{\dagger}(i\tilde{H}_{n})^{-1}V_{+,n}$
+$-V_{+,n}^{\dagger}(-i\tilde{H}_{n}^{*})^{-1}V_{+,n}\otimes P_{g},$
+where we used the identity $\int dte^{i\tilde{H}t}=(i\tilde{H})^{-1}e^{i\tilde{H}t}$. In the original representation for density matrix, it corresponds to the non-Hermitian effective Hamiltonian
 $\mathcal{H}^{\rm eff}_{n}=-V_{+,n}^{\dagger}(\tilde{H}_{n})^{-1}V_{+,n}=-\frac{\Omega^{2}\tilde{\delta}}{\tilde{\Delta}\tilde{\delta}-g^{2}A_{n}^{2}}|gn\rangle\langle gn|.$ (45)
-By substituting Eqs. (36)-(38) into the second term of Eq. (43), we can obtain the effective superoperator for quantum jumps:
-[…]
+By substituting Eqs. (36)-(38) into the second term of Eq. (43), one can obtain the effective superoperators for quantum jump:
+$\mathcal{P}_{g}\mathcal{J}_{n}(t)\int_{0}^{t}dt_{1}\mathcal{V}_{n}(t_{1})\int_{0}^{t_{1}}dt_{2}\mathcal{V}_{n}(t_{2})\mathcal{P}_{g}$ (46)
 $=$ $L_{hg,n}(\tilde{H}_{n}^{*})^{-1}V_{+,n}\otimes L_{hg,n}(\tilde{H}_{n})^{-1}V_{+,n}$
@@
 By transforming back to the original representation, the second-order effective Lindblad operators turn out to be
-$L_{hg,n}^{\text{eff}}=$ $L_{hg,n}(\tilde{H}_{n})^{-1}V_{+,n}=\frac{\sqrt{\gamma_{hg}}\Omega\tilde{\delta}}{\tilde{\Delta}\tilde{\delta}-g^{2}A_{n}^{2}}|gn\rangle\langle gn|,$ (47)
-$L_{eg,n}^{\text{eff}}=$ $L_{eg,n}(\tilde{H}_{n})^{-1}V_{+,n}=\frac{\sqrt{\gamma_{eg}}\Omega gA_{n}}{g^{2}A_{n}^{2}-\tilde{\Delta}\tilde{\delta}}|g(n+1)\rangle\langle gn|.$ (48)
-Therefore according to Eq. (43), the equation of motion in the ground state subspaces reads
-$P_{g}\dot{\rho}_{n}(t)P_{g}=-i\left[\mathcal{H}^{\rm eff}_{n}\rho_{n}(t)-\rho_{n}(t)(\mathcal{H}^{\rm eff}_{n})^{\dagger}\right]$ (49)
+$L_{hg,n}^{\rm eff}=$ $L_{hg,n}(\tilde{H}_{n})^{-1}V_{+,n}=\frac{\sqrt{\gamma_{hg}}\Omega\tilde{\delta}}{\tilde{\Delta}\tilde{\delta}-g^{2}A_{n}^{2}}|gn\rangle\langle gn|,$ (47)
+$L_{eg,n}^{\rm eff}=$ $L_{eg,n}(\tilde{H}_{n})^{-1}V_{+,n}=\frac{\sqrt{\gamma_{eg}}\Omega gA_{n}}{g^{2}A_{n}^{2}-\tilde{\Delta}\tilde{\delta}}|g(n+1)\rangle\langle gn|.$ (48)
+Thus according to Eq. (43), the equation of motion in the ground-state subspaces reads
+$P_{g}\dot{\rho}_{n}(t)P_{g}\approx-i\left[\mathcal{H}^{\rm eff}_{n}\rho_{n}(t)-\rho_{n}(t)(\mathcal{H}^{\rm eff}_{n})^{\dagger}\right]$ (49)
 $+$ $L_{hg,n}^{\rm eff}\rho(t)(L_{hg,n}^{\rm eff})^{\dagger}+L_{eg,n}^{\rm eff}\rho(t)(L_{eg,n}^{\rm eff})^{\dagger}.$
 It can be rewritten as a standard master equation
-$P_{g}\dot{\rho}_{n}(t)P_{g}=$ $-i\left[H^{\text{eff}}_{n},\rho_{n}(t)\right]+\mathcal{L}[L_{hg,n}^{\rm eff}]\rho_{n}(t)$ (50)
+$P_{g}\dot{\rho}_{n}(t)P_{g}\approx$ $-i\left[H^{\rm eff}_{n},\rho_{n}(t)\right]+\mathcal{L}[L_{hg,n}^{\rm eff}]\rho_{n}(t)$ (50)
 $+\mathcal{L}[L_{eg,n}^{\rm eff}]\rho_{n}(t)$
@@
 It is the Hermitian part of the operator
-$h_{n}^{\rm eff}=\mathcal{H}^{\rm eff}_{n}+\frac{i}{2}\sum_{k=hg,eg}(L^{\rm eff}_{k,n})^{\dagger}L^{\rm eff}_{k,n}$ (52)
-$=$ $\mathcal{H}^{\rm eff}_{n}+\frac{i}{2}V_{+,n}^{\dagger}(\tilde{H}_{n}^{\dagger})^{-1}\sum_{k=hg,eg,he}L_{k,n}^{\dagger}L_{k,n}\tilde{H}_{n}^{-1}V_{+,n}$
-$-\frac{i}{2}V_{+,n}^{\dagger}(\tilde{H}_{n}^{\dagger})^{-1}L_{he,n}^{\dagger}L_{he,n}\tilde{H}_{n}^{-1}V_{+,n}$
+$\mathcal{H}^{\rm eff}_{n}+\frac{i}{2}\sum_{k=hg,eg}\left(L^{\rm eff}_{k,n}\right)^{\dagger}L^{\rm eff}_{k,n}$ (52)
 $=$ $\mathcal{H}^{\rm eff}_{n}+\frac{1}{2}V_{+,n}^{\dagger}(\tilde{H}_{n}^{\dagger})^{-1}\left(\tilde{H}_{n}^{\dagger}-\tilde{H}_{n}\right)\tilde{H}_{n}^{-1}V_{+,n}$
@@
 $=$ $\mathcal{H}^{\rm eff}_{n}-\frac{1}{2}V_{+,n}^{\dagger}\left[(\tilde{H}_{n}^{\dagger})^{-1}-\tilde{H}_{n}^{-1}\right]V_{+,n}-\frac{i}{2}(L^{\rm eff}_{he,n})^{\dagger}L^{\rm eff}_{he,n}$
-$=$ $H^{\rm eff}_{n}-\frac{i}{2}(L^{\rm eff}_{he,n})^{\dagger}L^{\rm eff}_{he,n},$
-where $L_{he,n}=\sqrt{\gamma_{he}}|en\rangle\langle hn|$.
+$=$ $H^{\rm eff}_{n}-\frac{i}{2}(L^{\rm eff}_{he,n})^{\dagger}L^{\rm eff}_{he,n}.$
 The non-Hermitian part represents the leakage from $|gn\rangle$ to $|en\rangle$ with the jump operator $L^{\rm eff}_{he,n}=L_{he,n}\tilde{H}_{n}^{-1}V_{+,n}$.
 It is negligible comparing to the spontaneous decay $\gamma_{eg}$ under the condition $\Delta,\gamma_{hg}\gg\gamma_{he},\Omega$.
-For the uppermost ground-state subspace spanned by $\{|gN\rangle,|hN\rangle\}$, the non-Hermitian Hamiltonian reads
+For the uppermost subspace spanned by $\{|gN\rangle,|hN\rangle\}$, the non-Hermitian Hamiltonian reads
 $\tilde{H}_{N}=\tilde{\Delta}|hN\rangle\langle hN|.$ (53)
 Following the same procedure, the effective Lindblad operator and the effective Hamiltonian turn out to be
-$L^{\rm eff}_{hg,N}=$ $L_{hg,N}\tilde{H}_{N}V_{+,N}=\frac{\sqrt{\gamma_{hg}}\Omega}{\tilde{\Delta}}|gN\rangle\langle gN|,$ (54)
+$L^{\rm eff}_{hg,N}=$ $L_{hg,N}(\tilde{H}_{N})^{-1}V_{+,N}=\frac{\sqrt{\gamma_{hg}}\Omega}{\tilde{\Delta}}|gN\rangle\langle gN|,$ (54)
 $H^{\rm eff}_{N}=$ $-\Omega^{2}{\rm Re}\left(\frac{1}{\tilde{\Delta}}\right)|gN\rangle\langle gN|,$ (55)
 respectively. Up to this point, we have established the effective master equation describing the main unidirectional charging mechanism in Fig. 2(a).
-The leading-order correction induced by the decay $\gamma_{he}$ is described in Fig.
+The leading-order discharging induced by the decay $\gamma_{he}$ is described in Fig.
 2(b).
-The process $|gn\rangle\rightarrow|hn\rangle\rightarrow|en\rangle\rightarrow|gn\rangle$ gives rise to a pure dephasing effect on $|gn\rangle$, which is associated with a higher order term $\mathbf{L}^{(4)}_{{\rm eff},n}$ in Eq.
+The three-step transition $|gn\rangle\rightarrow|hn\rangle\rightarrow|en\rangle\rightarrow|gn\rangle$ gives rise to a pure dephasing effect on $|gn\rangle$.
+It is associated with a higher order term $\mathbf{L}^{(4)}_{{\rm eff},n}$ in Eq.
 (42)
-[…]
-×$\times\int_{0}^{t_{1}}dt_{2}\mathcal{V}_{n}(t_{2})\int_{0}^{t_{2}}dt_{3}\mathcal{V}_{n}(t_{3})\mathcal{P}_{g},$
+$J^{\text{eff}*}_{eg,n}\otimes J^{\rm eff}_{eg,n}=$ $\mathcal{P}_{g}\tilde{L}^{*}_{eg,n-1}(t)\otimes\tilde{L}_{eg,n-1}(t)\int_{0}^{t}dt_{1}\mathcal{J}_{he,n}(t_{1})$ (56)
+$\times\int_{0}^{t_{1}}dt_{2}\mathcal{V}_{n}(t_{2})\int_{0}^{t_{2}}dt_{3}\mathcal{V}_{n}(t_{3})\mathcal{P}_{g},$
 where $\mathcal{J}_{he,n}(t)$ describes the decay $\gamma_{he}$:
 $\mathcal{J}_{he,n}(t)=e^{-i\tilde{H}_{n-1}^{*}t}L_{he,n}e^{i\tilde{H}_{n}^{*}t}\otimes e^{i\tilde{H}_{n-1}t}L_{he,n}e^{-i\tilde{H}_{n}t}$ (57)
-with $L_{he,n}=\sqrt{\gamma_{he}}|en\rangle\langle hn|$. The dephasing rate turns out to be
+with $L_{he,n}=\sqrt{\gamma_{he}}|en\rangle\langle hn|$. The dephasing rate turns out to satisfy
 $\frac{|\langle gn|J^{\rm eff}_{eg,n}|gn\rangle|^{2}}{|\langle gn|L^{\rm eff}_{hg,n}|gn\rangle|^{2}}=\frac{\gamma_{eg}\gamma_{he}}{\gamma_{hg}}\mathbf{H}^{-1}_{en,en},$ (58)
@@
 $\mathbf{H}^{-1}_{en,en}=\langle en|\otimes\langle en|$ (59)
-×$\times\left(-i\tilde{H}^{*}_{n-1}\otimes\mathbb{I}+\mathbb{I}\otimes i\tilde{H}_{n-1}\right)^{-1}|en\rangle\otimes|en\rangle$
+$\times\left(-i\tilde{H}^{*}_{n-1}\otimes\mathbb{I}+\mathbb{I}\otimes i\tilde{H}_{n-1}\right)^{-1}|en\rangle\otimes|en\rangle,$
 using the identity $e^{-i\tilde{H}^{*}t}\otimes e^{i\tilde{H}t}=\exp(-i\tilde{H}^{*}t\otimes\mathbb{I}+\mathbb{I}\otimes i\tilde{H}t)$.
-Similarly, the process $|gn\rangle\rightarrow|hn\rangle\rightarrow|en\rangle\rightarrow|h(n-1)\rangle\rightarrow|g(n-1)\rangle$ in Fig.
+Similarly, the four-step transition $|gn\rangle\rightarrow|hn\rangle\rightarrow|en\rangle\rightarrow|h(n-1)\rangle\rightarrow|g(n-1)\rangle$ in Fig.
 2(b) gives rise to an effective transition from $|gn\rangle$ to $|g(n-1)\rangle$, the relevant jump superoperator is given by
 $J^{\rm eff*}_{hg,n}\otimes J^{\rm eff}_{hg,n}=\mathcal{P}_{g}\tilde{L}^{*}_{hg,n-1}(t)\otimes\tilde{L}_{hg,n-1}(t)$ (60)
-×$\times\int_{0}^{t}dt_{1}\mathcal{J}_{he,n}(t_{1})\int_{0}^{t_{1}}dt_{2}\mathcal{V}_{n}(t_{2})\int_{0}^{t_{2}}dt_{3}\mathcal{V}_{n}(t_{3})\mathcal{P}_{g}.$
+$\times\int_{0}^{t}dt_{1}\mathcal{J}_{he,n}(t_{1})\int_{0}^{t_{1}}dt_{2}\mathcal{V}_{n}(t_{2})\int_{0}^{t_{2}}dt_{3}\mathcal{V}_{n}(t_{3})\mathcal{P}_{g}.$
 The ratio of the effective decay rate from $|gn\rangle$ to $|g(n-1)\rangle$ in Fig. 2(b) to that from $|gn\rangle$ to $|g(n+1)\rangle$ in Fig. 2(a) is
@@
 $\mathbf{H}^{-1}_{h(n-1),en}=\langle h(n-1)|\otimes\langle h(n-1)|$ (62)
-×$\times\left(-i\tilde{H}^{*}_{n-1}\otimes\mathbb{I}+\mathbb{I}\otimes i\tilde{H}_{n-1}\right)^{-1}|en\rangle\otimes|en\rangle.$
-Under the parametric settings that $\gamma_{eg}\approx\gamma_{he}\ll\gamma_{hg}$, it is important to observe that the dephasing rate given by Eq.
-(58) is negligible in comparison to that in Eq.
-(47).
-Similarly, the effective transition rate from $|gn\rangle$ to $|g(n-1)\rangle$ given by Eq.
-(61) is negligible in comparison to that from $|gn\rangle$ to $|g(n+1)\rangle$ in Eq.
-(48) under the condition $\gamma_{eg}\approx\gamma_{he}\ll\gamma_{hg}$ and $\Delta\gg\delta$ due to our optimization condition $g^{2}A_{n}^{2}\approx|\tilde{\Delta}\tilde{\delta}|$.
+$\times\left(-i\tilde{H}^{*}_{n-1}\otimes\mathbb{I}+\mathbb{I}\otimes i\tilde{H}_{n-1}\right)^{-1}|en\rangle\otimes|en\rangle.$
+[…]